### v1

# Quasirandom and quasisimple groups

Abstract.
Fix $\varepsilon>0$. We say that a finite group $G$ is $\varepsilon$-quasirandom if every nontrivial irreducible complex representation of $G$ has degree at least $|G|^{\varepsilon}$. In this paper, we give a structure theorem for large $\varepsilon$-quasirandom groups, and we completely classify the $\frac{1}{5}$-quasirandom groups.
Key words and phrases:
Quasirandom groups, quasisimple groups.
2020 Mathematics Subject Classification:
Primary: 20D99, 20C99.
1. Introduction
Informally, a finite group is called quasirandom if its nontrivial complex irreducible representations have “large” degree, where the precise meaning of “large” depends on context. The terminology was introduced explicitly in [19], where the name arises because dense Cayley graphs in quasirandom groups are quasirandom graphs in the sense of Chung, Graham, and Wilson [6]. This property had already been used implicitly to construct expander graphs in [23, 9]. After Gowers’ groundbreaking paper, quasirandom groups have been at the center of spectacular development. Bourgain and Gamburd [3] used the quasirandomness of $\mathrm{SL}_{2}(p)$ to show that random Cayley graphs on these groups are expanders with high probability. Their work has inspired an impressive series of articles, culminating in [4]. In a more subtle way, quasirandomness played a key role in Helfgott’s work [11] and generalizations, most notably [5, 21, 8]. Tao dedicated an entire chapter of his book [26] to quasirandom groups.
As said, being quasirandom is a quantitative property. For applications to expansion in Cayley graphs, one needs that every nontrivial irreducible representation has degree at least $|G|^{\varepsilon}$, for some constant $\varepsilon>0$ independent of the size of $G$. This motivates the following definition:
Definition.
Let $\varepsilon>0$. A finite group $G$ is $\varepsilon$-quasirandom if every nontrivial irreducible complex representation of $G$ has degree at least $|G|^{\varepsilon}$.
It is well known that the finite simple groups of Lie type are $\varepsilon$-quasirandom for some $\varepsilon>0$ which depends only on the rank [14], a fact that is exploited in all the papers we mentioned above. A more general discussion on quasirandomness is taken by Babai, Nikolov and Pyber [2], and Nikolov and Pyber [18]. The celebrated Gower’s trick [18] says that, if $G$ is $\varepsilon$-quasirandom and $S\subseteq G$ is a large subset, namely $|S|\geq|G|^{1-\frac{\varepsilon}{3}}$, then $S^{3}=G$. This fact, as well as other natural invariants of $G$, can be used to characterize quasirandomness. Such equivalences look particularly nice when expressed in the exponential regime (see Theorem 2.6 below).
The algebraic structure of quasirandom groups remains unexplored. The present article was born from our need to answer the question: what, concretely, is a quasirandom group? In the first main result, we prove that large $\varepsilon$-quasirandom groups with a trivial solvable radical are direct products of nonabelian simple groups. We say a little more.
Theorem A.
Fix $\varepsilon>0$, and let $G$ be an $\varepsilon$-quasirandom group. If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/Sol(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
For the second main result, we study the groups that are $\varepsilon$-quasirandom with a relatively large $\varepsilon$. It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the most quasirandom family of groups [24, 29], but not much is known. Not only this is actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
Theorem B.
Let $G$ be a $\frac{1}{5}$-quasirandom group. Then one of the following holds:
  1. $(a)$
    $G\cong(\mathbb{F}_{3})^{3}\rtimes\mathrm{SL}_{3}(3)$;
  2. $(b)$
    $G$ is a perfect central extension of $(\mathbb{F}_{2^{f}})^{4}\rtimes\mathrm{Sp}_{4}(2^{f})$, $f\geq 6$;
  3. $(c)$
    $G$ is quasisimple.
The irreducible representations of the quasisimple groups are known [27], thus it is an easy (although laborious) task to recover the full list of the $\frac{1}{5}$-quasirandom groups from Theorem B. An interesting consequence is the following:
Corollary C.
Every $\frac{3}{14}$-quasirandom group is quasisimple.
The constant $\frac{3}{14}$ is best possible, precisely because of the family in Theorem Bb. On the other hand, the value $\frac{1}{5}$ is the result of a compromise, as with our machinery and more work one can obtain a deeper classification. There are some reasons to prefer $\frac{1}{5}$: first, it is small enough to provide Corollary C. Second, it makes the theorem sharp in some sense, because there exist infinitely many nonquasisimple groups which are $(\frac{1}{5}-\varepsilon)$-quasirandom, for any $\varepsilon>0$ (see Section 3). It also makes the proof a little cleaner, because every $\frac{1}{5}$-quasirandom group has a unique maximal normal subgroup (Lemma 4.7). We point out that similar methods are useful to study quasirandomnes at other regimes different from the exponential, which could be of interest as well. The shallow version of Theorem B is the following:
Corollary D.
The only $\frac{1}{3}$-quasirandom groups are the sporadic simple groups $J_{1}$ and $O^{\prime}N$.
We will actually prove Corollary D as a “baby” Theorem B. The philosophy behind the proofs is that quasirandom groups are pretty rigid, and lend themselves very little to nontrivial extensions, especially the noncentral ones. To make this idea really say something, we use various properties of the finite simple groups and their representations, both complex and modular, linear and projective.
The paper is structured as follows. In Section 2 we introduce the (exponential) quasirandom degree $\bm{\varepsilon}(G)$ of a finite group, and give basic results about quasisimple groups, modular and projective representations. In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem Bb. We prove the main theorems in Section 4. Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups: we collect these data in the Appendix.
2. Preliminaries
2.1. Minimal degree of a nontrivial representation
In this paper, every group is finite. For a nontrivial finite group $G$, we denote by $\mathbf{D}_{0}(G)$ the minimal degree of a nontrivial irreducible complex representation of $G$, that is,
$\mathbf{D}_{0}(G)\>:=\>\min\left\{\dim(\pi)\mid\pi\in\mathrm{Irr}(G)\setminus\{1\}\right\}.$
We start with the following, which appears as [26, Ex. 1.3.2].
Lemma 2.1.
Let $N\lhd G$ be a nontrivial proper normal subgroup. Then
$\mathbf{D}_{0}(G/N)\,\geq\,\mathbf{D}_{0}(G)\,.$
If the previous inequality is strict, then
$\mathbf{D}_{0}(G)\>\geq\>\mathbf{D}_{0}(N).$
Proof.
Each nontrivial irreducible representation of $G/N$ provides a nontrivial representation of $G$. The first inequality follows. For the second one, suppose there exists $\pi\in\mathrm{Irr}(G)$ such that $\dim(\pi)=\mathbf{D}_{0}(G)<\mathbf{D}_{0}(G/N)$. Observe that the restriction $\pi_{\downarrow_{N}}$ of $\pi$ to $N$ is nontrivial: otherwise $N\subseteq Ker(\pi)$, and $\pi$ would induce a nontrivial irreducible representation of $G/N$ whose degree is $\dim(\pi)$, against the minimality of $\mathbf{D}_{0}(G/N)$. Therefore, $\pi_{\downarrow N}$ contains a nontrivial irreducible representation of $N$, and so $\dim(\pi)=\dim(\pi_{\downarrow_{N}})\geq\mathbf{D}_{0}(N)$. ∎
Lemma 2.2.
If $G$ acts nontrivially on a finite set $\Omega$, then $\mathbf{D}_{0}(G)\leq|\Omega|-1$.
Proof.
Recall that $\mathrm{Sym}(n)$ has a natural embedding into $\mathrm{GL}_{n}$, and that any permutation representation contains the trivial representation, because the diagonal subspace is fixed by every element of $G$. ∎
We can obtain some useful bounds by considering the natural actions of $G$.
Corollary 2.3.
The following inequalities hold:
  1. $(i)$
    if $H<G$ is a proper subgroup, then
    $|G:H|\,\geq\,\mathbf{D}_{0}(G/\mathrm{Core}_{G}(H))+1\,;$
  2. $(ii)$
    if $N\lhd G$ is a noncentral normal subgroup, then
    $|N|\,\geq\,\mathbf{D}_{0}(G/\mathbf{C}_{G}(N))+2\,;$
  3. $(iii)$
    if $C\subseteq G$ is a nontrivial conjugacy class, then
    $|C|\,\geq\,\mathbf{D}_{0}(G)+1\,.$
Proof.
Use Lemma 2.2 on: the action of $G/\mathrm{Core}_{G}(H)$ on the right cosets of $H$ by multiplication; the action of $G/\mathbf{C}_{G}(N)$ on $N\setminus\{1\}$ by conjugation; the action of $G$ on $C$ by conjugation. ∎
We would like to mention some connections between $\mathbf{D}_{0}(G)$ and other important parameters associated to a finite group: the number of conjugacy classes $\mathbf{k}(G)$, the representation growth $\mathbf{r}_{n}(G)$, and the minimal index of a proper subgroup, namely $\bm{\iota}(G):=\min_{H<G}|G:H|$.
Remark 2.4.
By standard results in representation theory, we can write
$|G|=\sum_{\pi\in\mathrm{Irr}(G)}\dim(\pi)^{2}\geq 1+(\mathbf{k}(G)-1)\mathbf{D}_{0}(G)^{2}.$
It follows that, if $G$ is an $\varepsilon$-quasirandom group, then $\mathbf{k}(G)\leq|G|^{1-2\varepsilon}$. Moreover, the number $\mathbf{r}_{n}(G)$ of the inequivalent irreducible representations of $G$ of degree $n$ is polynomially bounded, i.e. $\mathbf{r}_{n}(G)\leq n^{\varepsilon^{-1}-2}$ for all $n$ (see [22, Lemma 25]).
Remark 2.5.
Using the Classification of the Finite Simple Groups, Nikolov and Pyber [18] have proved that there is an absolute constant $c$ such that
$\bm{\iota}(G)\>\leq\>c\>\mathbf{D}_{0}(G)^{2}$
for every finite perfect group $G$. A converse inequality, $\mathbf{D}_{0}(G)<\bm{\iota}(G)$, follows from the previous lemmas. Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$). In particular, $\mathbf{D}_{0}(G)<|G|^{0.5}$. As consequences of Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
2.2. Quasirandom degree
Let $G$ be a nontrivial finite group. We introduce the (exponential) quasirandom degree of $G$ by
$\bm{\varepsilon}(G)\>:=\>\log_{|G|}\mathbf{D}_{0}(G).$
By definition, $\mathbf{D}_{0}(G)=|G|^{\bm{\varepsilon}(G)}$, and $\bm{\varepsilon}(G)$ is the maximum $\varepsilon\in[0,1)$ such that $G$ is an $\varepsilon$-quasirandom group. Observe that $\bm{\varepsilon}(G)>0$ if and only if $G$ is perfect. We now report the theorem of Nikolov and Pyber mentioned in the Introduction:
Theorem 2.6 (Theorem 3.1 in [18]).
Let $G$ be a finite perfect group. The following statements are roughly equivalent, in the sense that if one statement holds for a small constant $\varepsilon>0$, then all others hold with constants that are a positive fraction of $\varepsilon$:
  • $G$ is $\varepsilon$-quasirandom;
  • every proper subgroup of $G$ has index at least $|G|^{\varepsilon}$;
  • for every subset $S$ of size at least $|G|^{1-\varepsilon}$ we have $S^{3}=G$;
  • every product-free subset of $G$ has size at most $|G|^{1-\varepsilon}$.
Let $N\lhd G$ be a proper normal subgroup. From an easy computation, we write
$\bm{\varepsilon}(G)\>=\>\bm{\varepsilon}(G/N)\cdot\frac{\log\mathbf{D}_{0}(G)}{\log\mathbf{D}_{0}(G/N)}\cdot\frac{\log|G/N|}{\log|G|}.$
(We remark here that, when we consider a ratio of logarithms, the base can be chosen arbitrarily.) It is immediate from the equality above and Lemma 2.1 that $\bm{\varepsilon}(G/N)\geq\bm{\varepsilon}(G)$, with equality if and only if $N$ is trivial. We will use this fact throughout the article without mentioning it.
We now give a simple but very useful inequality. Let $N\lhd G$ be a proper normal subgroup. By the definition of quasirandom degree, and Lemma 2.1,
$|G|\>=\>\mathbf{D}_{0}(G)^{1/\bm{\varepsilon}(G)}\>\leq\>\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}.$
Hence, substituting $|G|=|G/N||N|$, we obtain
(2.1) $|N|\>\leq\>\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.$
What makes (2.1) so useful is that in many occasions we barely know that $\bm{\varepsilon}(G)$ is large, but we have a good knowledge of $G/N$. The next lemma provides a gap between $\bm{\varepsilon}(G)$ and $\bm{\varepsilon}(G/N)$ when $N$ is abelian but noncentral, and it greatly simplifies the proof of Theorem B.
Lemma 2.7.
Let $N\lhd G$ be a nontrivial noncentral abelian normal subgroup. Then
$\bm{\varepsilon}(G/N)\>>\>\frac{\bm{\varepsilon}(G)}{1-\bm{\varepsilon}(G)}.$
Proof.
Since $N$ is abelian, we have $N\subseteq\mathbf{C}_{G}(N)\neq G$. Combining Lemma 2.1, Corollary 2.3(ii), and (2.1), we write
$\begin{split}\mathbf{D}_{0}(G/N)&\leq\mathbf{D}_{0}(G/\mathbf{C}_{G}(N))\\ &<|N|\\ &\leq\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.\end{split}\,$
Writing $\mathbf{D}_{0}(G/N)=|G/N|^{\bm{\varepsilon}(G/N)}$, we obtain
$|G/N|^{\bm{\varepsilon}(G/N)}<|G/N|^{\frac{\bm{\varepsilon}(G/N)}{\bm{\varepsilon}(G)}-1},$
and the result is proven by comparing the exponents. ∎
Lemma 2.7 may fail if $N$ is central: for instance, we have
$\bm{\varepsilon}(\mathrm{SL}_{2}(7))\approx 0.1889,\quad\bm{\varepsilon}(\mathrm{PSL}_{2}(7))\approx 0.2144,$
and
$\bm{\varepsilon}(\mathrm{PSL}_{2}(7))\><\>\frac{\bm{\varepsilon}(\mathrm{SL}_{2}(7))}{1-\bm{\varepsilon}(\mathrm{SL}_{2}(7))}.$
Such a phenomenon suggests that quasisimple groups are strong candidates looking for highly quasirandom groups, because $\bm{\varepsilon}(G)$ can be very close to $\bm{\varepsilon}(G/\mathbf{Z}(G))$.
We now estimate the quasirandom degree of a direct product.
Lemma 2.8.
Let $A$ and $B$ be two finite groups, and suppose that $\bm{\varepsilon}(A)\leq\bm{\varepsilon}(B)$. Then
$\frac{\bm{\varepsilon}(A)}{2}\>\leq\>\bm{\varepsilon}(A\times B)\>\leq\>\frac{\bm{\varepsilon}(B)}{2}.$
Proof.
We have $\mathbf{D}_{0}(A\times B)=\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}$, so that we can write
$\begin{split}\bm{\varepsilon}(A\times B)&=\frac{\log\left(\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}\right)}{\log(|A||B|)}\\ &=\frac{\log\left(\min\{|A|^{\bm{\varepsilon}(A)},|B|^{\bm{\varepsilon}(B)}\}\right)}{\log|A|+\log|B|}\\ &\leq\bm{\varepsilon}(B)\cdot\frac{\min\{\log|A|,\log|B|\}}{\log|A|+\log|B|}\\ &\leq\frac{\bm{\varepsilon}(B)}{2}.\end{split}\,$
The proof of the other inequality is similar. ∎
Corollary 2.9.
Let $(G_{i})_{i=1}^{n}$ be finite groups and suppose $\bm{\varepsilon}(G_{1})\leq...\leq\bm{\varepsilon}(G_{n})$. Then $\bm{\varepsilon}(A_{1})/n\leq\bm{\varepsilon}(\prod_{i=1}^{n}A_{i})\leq\bm{\varepsilon}(A_{n})/n$.
2.3. Quasisimple groups
We recall that a central extension of a group $Q$ is a short exact sequence
[…]
where $N\subseteq\mathbf{Z}(G)$. If $G$ is perfect, then $G$ is a perfect central extension. A finite group is quasisimple if it is a perfect central extension of a nonabelian finite simple group, that is, $G$ is perfect and $G/\mathbf{Z}(G)$ is simple.
It is a key property of perfect groups that, if $G$ is perfect, then there is a unique universal central perfect extension, which we denote by $\widetilde{G}$. This extension is obtained by choosing $N$ isomorphic to the Schur multiplier of $G$, that is, the second homology group $H^{2}(G,\mathbb{Z})$. For more details on this, we refer to [1, Sec. 33].
The next two lemmas are essentially well known. We write down Lemma 2.11 below, which is a little more general, and will be useful later.
Lemma 2.10.
Let $G$ be a quasisimple group, and let $N\lhd G$ be a proper normal subgroup. Then $N\subseteq\mathbf{Z}(G)$, and $G/N$ is again quasisimple. Moreover, a perfect central extension of a quasisimple group is quasisimple.
Proof.
See [1, (31.2) and (33.5)]. ∎
Lemma 2.11.
Let $G$ be a finite perfect group. Let $N\subseteq\mathbf{Z}(G)$, and $N\lhd N_{0}\lhd G$ such that $N_{0}/N\subseteq\mathbf{Z}(G/N)$. Then $N_{0}\subseteq\mathbf{Z}(G)$.
Proof.
Let $Z=\mathbf{Z}(G)$. We first show that $\mathbf{Z}(G/N)=Z/N$. It is clear that $Z/N\subseteq\mathbf{Z}(G/N)$. On the other hand, $\mathbf{Z}(G/N)$ is certainly contained in the projection of $\mathbf{Z}(G/Z)$ in $G/N$. By Grün’s Lemma [10] we have $\mathbf{Z}(G/Z)=1$, so $\mathbf{Z}(G/N)\subseteq Z/N$.
Now from the hypotheses we have $N_{0}/N\subseteq Z/N$, i.e. $N_{0}\subseteq Z$. ∎
We conclude this subsection with the following observation, which is useful for an inductive argument.
Lemma 2.12.
Let $G$ be a nonquasisimple group in which all the proper quotients are quasisimple. Then either $G$ has a unique minimal normal subgroup, or $G$ is the direct product of two nonabelian finite simple groups.
Proof.
We first notice that either $G$ is a group of prime order, or it is perfect. In the first case the statement is trivial, thus we can suppose that $G$ is perfect.
Let us suppose that $N_{1}$ and $N_{2}$ are two distinct minimal normal subgroups of $G$, so that $N_{1}\cap N_{2}=1$. Let $H=N_{1}N_{2}$, and note that $G/N_{i}$ is quasisimple for $i\in\{1,2\}$. If $G=H$, then $G$ is the direct product of finite simple factors. In particular, it is easy to see that these factors must be precisely two and nonabelian.
Aiming for a contradiction, suppose $H\neq G$. Since all the proper normal subgroups of a quasisimple group are central, $H/N_{i}\subseteq\mathbf{Z}(G/N_{i})$ for both $i\in\{1,2\}$. Let $h\in H$ and $g\in G$. For each $i\in\{1,2\}$, we have
$[h,g]N_{i}=[hN_{i},gN_{i}]\subseteq N_{i}.$
It follows that $[h,g]\in N_{1}\cap N_{2}=1$, and since $h\in H$ is arbitrary, we conclude $H\subseteq\mathbf{Z}(G)$. Finally, by Lemma 2.10, $G$ itself is quasisimple, against our initial assumption. ∎
2.4. Linear and projective representations
So far, we have only dealt with complex linear representations, but modular and projective representations will be a key tool in the proof of the main theorems. For a prime $r\geq 2$, we use $\mathbf{D}_{r}(G)$ to denote the minimal degree of a nontrivial irreducible modular representation of $G$ over a field of characteristic $r$. We recall that a projective representation of $G$ is a representation of a perfect central extension of $G$. Equivalently, it is a group homomorphism
$\rho:\>G\to\mathrm{PGL}(V).$
The degree of the representation is the dimension of $V$. As for linear representations, we denote by $\mathbf{R}_{0}(G)$ the minimal degree of a nontrivial projective complex representation, and by $\mathbf{R}_{r}(G)$ the minimal degree of a nontrivial projective representation over a field of characteristic $r$. We remark that, if $G$ is finite and perfect, and $\widetilde{G}$ is its universal perfect central extension, then
$\mathbf{R}_{0}(G)=\mathbf{D}_{0}(\widetilde{G})\leq\mathbf{D}_{0}(G),\quad\hbox{and}\quad\mathbf{R}_{r}(G)=\mathbf{D}_{r}(\widetilde{G})\leq\mathbf{D}_{r}(G).$
In the remainder of this subsection, we focus on the projective representations of finite simple groups of Lie type. We recall that, if $G$ is a group of Lie type defined over a field of characteristic $p$, then a modular representation of $G$ is in natural characteristic when the characteristic of the underlying vector space is $p$, while it is in cross-characteristic otherwise.
Lemma 2.13.
Let $L$ be a finite simple group of Lie type over $\mathbb{F}_{q}$, $q=p^{f}$. If $V$ is a vector space in natural characteristic $p$ on which $\widetilde{L}$ acts nontrivially, then
$|V|\>\geq\>q^{\mathbf{R}_{p}(L)}.$
Proof.
It follows from Propositions 5.3.1, 5.4.6, and Remark 5.4.7 of [13]. ∎
For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom simple groups. The infinite families of $\frac{1}{4}$-quasirandom groups are
$\mathrm{PSL}_{d}(q)\hbox{ with }d\in\{2,3\},\quad\mathrm{Sp}_{4}(q)\hbox{ with }q\hbox{ even},$
$G_{2}(q)\hbox{ with }q=3(\mathrm{mod}\,6),\quad^{2}B_{2}(q^{2}),\quad^{2}G_{2}(q^{2})\,.$
The relevant informations can be recovered from [14, 25] for classical groups, and from [15, Sec. 4] for exceptional groups. We collect these data in Table 1.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\mathbf{D}_{0}(L)$ $\mathbf{R}_{0}(L)$ $\mathbf{R}_{r}(L)$, $r\neq p$ $\mathbf{R}_{p}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{2}(q)$ $q$ even $q-1$ $q-1$ $q-1$ $2$
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{3}(q)$ $q\neq 2,4$ $q(q+1)$ $q(q+1)$ $q^{2}+q-1$ $3$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{4}(q)$ $q$ even $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $4$
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $4$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $\phi_{12}(q)$ $\phi_{12}(q)$ $\phi_{12}(q)-\delta_{2}$ $7$
Table 1. The $\frac{1}{4}$-quasirandom simple groups of Lie type.
We denote by $\phi_{\ell}(q)$ the $\ell$-th cyclotomic polynomial in the variable $q$. The value of $\delta_{1}$ is $1$ if $r=2$, $q=3^{f}$, $f\geq 2$, or $r=3$, $3$ divides $q-1$, and it is 00 otherwise. The value of $\delta_{2}$ is $1$ if $r=2$, and 00 otherwise.
Remark 2.14.
We observe that, if $L$ is a $\frac{1}{4}$-quasirandom simple group of Lie type, in cross-characteristic we have
$\mathbf{R}_{r}(L)\>\geq\>\mathbf{D}_{0}(L)-1.$
Moreover, we point out that the comparison between $\mathbf{R}_{0}(L)$ and $\mathbf{R}_{r}(L)$ can be explicitly found, for all simple groups of Lie type, in [27, Sec. 3].
3. Quasirandom groups of affine type
In this short section, we provide some highly quasirandom nonquasisimple groups. The key tool is Lemma 3.2 below. Before giving the results, we revise some notions from the representation theory of abelian groups. Let $V$ be a finite vector space, and let $V^{*}$ be its dual space. We remark that $V^{*}\cong Hom(V,\mathbb{C}^{*})$. If $\phi:V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}:V^{*}\rightarrow V^{*}$ is defined by
$\phi^{*}(f)\>:=\>\phi\circ f.$
Let $K$ be a finite group. If $\rho:K\rightarrow\mathrm{GL}(V)$ is a representation of $K$, then the dual representation $\rho^{*}:K\rightarrow\mathrm{GL}(V^{*})$ is defined by
$\rho^{*}(g)\>:=\>\rho(g^{-1})^{*}$
for every $g\in K$.
Lemma 3.1.
Let $V$ be a finite vector space, and $K\leqslant\mathrm{GL}(V)$. Then the number of orbits of $K$ on $V$ equals the number of orbits of $K$ on $V^{*}$.
Proof.
Let $V=(\mathbb{F}_{q})^{n}$ and $g\in K$. The number of $g$-fixed points in $V$ is $q^{\dim(Ker(g-I))}$, while the number of $g$-fixed points in $V^{*}$ is $q^{\dim(Ker((g^{-1})^{T}-I))}$, where $g^{T}$ is the transpose of $g$. But $\dim(Ker(g-I))$ is the multiplicity of $1$ as an eigenvalue of $g$, so the two quantities above are equal. Since $g$ is arbitrary, the proof follows from Burnside’s lemma. ∎
Lemma 3.2.
Let $G=V\rtimes K$ be a $2$-transitive affine group. Then
$\mathbf{D}_{0}(G)\>=\>\mathbf{D}_{0}(K).$
Proof.
Since $K$ is a quotient of $G$, it is sufficient to show that $\mathbf{D}_{0}(G)\geq\mathbf{D}_{0}(K)$. We first notice that $K$ acts nontrivially on $V\setminus\{0\}$, so $\mathbf{D}_{0}(K)\leq|V|-2$ from Lemma 2.2. Moreover, we can identify $V^{*}$ with $\mathrm{Irr}(V)$. Let $\pi\in\mathrm{Irr}(G)\setminus\{1\}$. If $\pi_{\downarrow V}$ is the trivial representation of $V$, then $\pi\in\mathrm{Irr}(G/V)$ is an irreducible representation of $K$. In particular, $\dim(\pi)\geq\mathbf{D}_{0}(K)$. On the other hand, if $\pi_{\downarrow V}$ is a nontrivial representation of $V$, then, by Clifford theory, there is an irreducible character $\theta\in\mathrm{Irr}(V)\setminus\{1\}$ such that $\dim(\pi)$ is a multiple of $|K:I_{K}(\theta)|$, where $I_{K}(\theta)$ is the inertia subgroup of $\theta$. From Lemma 3.1, $K$ acts transitively on $\mathrm{Irr}(V)\setminus\{1\}$. By the orbit-stabilizer theorem, we have
$|K:I_{K}(\theta)|=|\theta^{K}|=|V|-1.$
From the discussion above, we obtain
$\dim(\pi)\geq|V|-1>\mathbf{D}_{0}(K).$
Since $\pi\in\mathrm{Irr}(G)\setminus\{1\}$ is arbitrary, we conclude that $\mathbf{D}_{0}(G)\geq\mathbf{D}_{0}(K)$ as desired. ∎
Example 3.3.
Let $q\geq 4$ range among the powers of $2$, and let $G_{q}=(\mathbb{F}_{q})^{4}\rtimes\mathrm{Sp}_{4}(q)$. Since $\mathrm{Sp}_{4}(q)$ is transitive on $(\mathbb{F}_{q})^{4}\setminus\{0\}$, from Lemma 3.2 we have
$\begin{split}\bm{\varepsilon}(G_{q})&=\frac{\log\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))}{\log|G_{q}|}\\ &=\frac{\log\left(\dfrac{q(q-1)^{2}}{2}\right)}{\log(q^{8}(q^{4}-1)(q^{2}-1))}.\end{split}\,$
A direct computation shows that this value is more than $\frac{1}{5}$ for $q\geq 64$, and it lies between $\frac{3}{14}-\frac{1}{\log_{2}q}$ and $\frac{3}{14}$, for all $q$. In particular, $\lim_{q\rightarrow+\infty}\bm{\varepsilon}(G_{q})=\left(\frac{3}{14}\right)^{-}$. Similarly, if $q$ is a prime power and $H_{q}=(\mathbb{F}_{q})^{2}\rtimes\mathrm{SL}_{2}(q)$, then from Lemma 3.2 we have
$\bm{\varepsilon}(H_{q})=\frac{\log\mathbf{D}_{0}(\mathrm{SL}_{2}(q))}{\log|H_{q}|}.$
A direct computation shows that $\bm{\varepsilon}(H_{q})$ lies between $5^{-1}-q^{-1}$ and $5^{-1}$ for all $q$, so in particular $\lim_{q\rightarrow+\infty}\bm{\varepsilon}(H_{q})=\left(\frac{1}{5}\right)^{-}$.
4. Proof of the main theorems
4.1. Proof of Theorem A
The main idea behind the proof of Theorem A is the following. Under the hypothesis that $Sol(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$. If this is not the case, then $G/F^{*}(G)$ has a nontrivial simple quotient $L$. We use the quasirandomness of $L$ to obtain an upper bound on $|G|$, while Bender’s theorem and other considerations provide a lower bound on $|F^{*}(G)|$. These inequalities fight each other, giving the desired contradiction. A similar argument will be used in a portion of the proof of Theorem B. Going into details, we start reporting the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
Theorem 4.1 (Schreier conjecture).
If $L$ is a finite simple group, then the group of the outer automorphisms $\mathrm{Out}(L)$ is solvable.
The following two observations concern sections of finite groups.
Remark 4.2.
Let $N\lhd G$, and suppose that a section $H_{0}/N_{0}\cong L$ of $G$ is simple. Then $N\subseteq N_{0}$, or $NN_{0}=H_{0}$. In the first case $L$ appears as a section of $G/N$, while in the second case $L$ appears as a section of $N$.
Lemma 4.3.
Let $\Omega$ be a finite set, and let $N\lhd H\leqslant\mathrm{Sym}(\Omega)$. If $H/N\cong L$ is simple, then $\mathrm{Sym}(\Omega)$ has a subgroup isomorphic to $L$.
Proof.
Let $|\Omega|=n$, $x\in\Omega$, and work by induction on $n$. Let $H_{x}$ be stabilizer of $x$. We claim that $NH_{x}/N$ is a corefree subgroup of $H/N$ of index at most $n$. First,
$|(H/N):(NH_{x}/N)|=|H:NH_{x}|\leq|H:H_{x}|\leq n.$
Now, if $NH_{x}$ is neither $N$ nor $H$, then the proof is finished by the fact that $H/N$ is simple. In the first case $|H/N|\leq n$, and the proof follows by the Cayley embedding theorem. If $NH_{x}=H$, then $H_{x}/H_{x}\cap N\cong H/N\cong L$, hence by induction $\mathrm{Sym}(\Omega\setminus\{x\})$ and so $\mathrm{Sym}(\Omega)$ has a subgroup isomorphic to $L$, as desired. ∎
In the following remark, we discuss the asymptotic behaviour of $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$, when $L$ is a finite simple group, and $|L|\rightarrow+\infty$.
Remark 4.4.
For $n\geq 6$,
$\mathbf{D}_{0}(\mathrm{Alt}(n))=n-1,\quad\hbox{and}\quad\bm{\varepsilon}(\mathrm{Alt}(n))=\frac{\log(n-1)}{\log(n!/2)}.$
In particular, for $L=\mathrm{Alt}(n)$, we have
$\mathbf{D}_{0}(L)\sim\frac{\log|L|}{\log\log|L|},\quad\hbox{and}\quad\bm{\varepsilon}(L)\sim\frac{\log\log|L|}{\log|L|}$
when $|L|\rightarrow+\infty$. These are the less quasirandom among the nonabelian finite simple groups. If $L=L_{d}(q)$ is a finite simple group of Lie type, then it is easy to check from the Appendix that
$q^{10^{-1}d}\>\leq\>\mathbf{D}_{0}(L_{d}(q))\>\leq\>q^{10d},$
and
[…]
In particular, if $L=L_{d}(q)$ and $q$ is fixed, then $\mathbf{D}_{0}(L)$ is comparable with $c^{\sqrt{\log|L|}}$, while $\bm{\varepsilon}(L)$ is comparable with $\frac{c}{\sqrt{\log|L|}}$.
Proof of Theorem A.
Fix $\varepsilon>0$. From Corollary 2.3(i), it is clear that if $|G|$ is large then $|G/Sol(G)|$ is large. Moreover, $G/Sol(G)$ is again $\varepsilon$-quasirandom. Up to replacing $G$ with $G/Sol(G)$, we can assume that $G$ has a trivial solvable radical. Now it is well known that the generalized Fitting subgroup $F=F^{*}(G)$ coincides with the socle of $G$, and it is the direct product of nonabelian simple groups $T_{i}$’s. Let $F=\prod_{i}(T_{i})^{n_{i}}$, where $n_{i}\geq 1$ for each $i$. Moreover, as $\mathbf{C}_{G}(F)\subseteq F$ [1, (31.13)], we have $\mathbf{C}_{G}(F)=\mathbf{Z}(F)=1$. It follows that $G$ can be embedded as a subgroup of
$\mathrm{Aut}(F)\>\cong\>\prod_{i}\mathrm{Aut}(T_{i})^{n_{i}}\rtimes\mathrm{Sym}(n_{i}).$
In particular, all the composition factors of $G/F$ lie as sections of one among $\mathrm{Out}(T_{i})$ or $\mathrm{Sym}(n_{i})$, for some $i$. We claim that $G=F$. Otherwise, it is easy to see that $G/F$ has a simple quotient $L$ which is $\varepsilon$-quasirandom. Of course, $|G|\leq\mathbf{D}_{0}(L)^{1/\varepsilon}$. So, if $|L|$ is bounded, then $|G|$ is bounded. It follows that $L=L_{d}(q)$ is a simple group of Lie type, $d$ is $\varepsilon$-bounded, while $q$ is unbounded. By the definition of quasirandom degree, we have
(4.1) $|G|\>\leq\>\mathbf{D}_{0}(L_{d}(q))^{1/\varepsilon}.$
Let $n=\max_{i}n_{i}$, so that $|F|\geq 60^{n}$. From Remark 4.2 and Theorem 4.1, $L_{d}(q)$ has to appear as a section of $\mathrm{Alt}(n)$. From Lemma 4.3, $\mathrm{Sym}(n)$ has a subgroup isomorphic to $L_{d}(q)$. From Lemma 2.2 we obtain $\mathbf{D}_{0}(L_{d}(q))\leq n$, and so
[…]
When $\mathbf{D}_{0}(L_{d}(q))$ is sufficiently large, this gives a contradiction to (4.1). But $\mathbf{D}_{0}(L_{d}(q))\geq|G|^{\varepsilon}$, so we obtain a contradiction when $|G|$ is sufficiently large.
We have proven that $G/Sol(G)=F^{*}(G/Sol(G))$, so $G/Sol(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/Sol(G)=\prod_{i}(T_{i})^{n_{i}}$. From Remark 4.4, the $T_{i}$’s are all simple groups of Lie type. Let $N=\sum_{i}n_{i}$. From Corollary 2.9, we have $\varepsilon\leq\mathbf{D}_{0}(G/Sol(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$. Moreover, each block $(T_{i})^{n_{i}}$ has to be $\varepsilon$-quasirandom as a quotient of $G$. From Corollary 2.9 and Remark 4.4, we have
$\varepsilon\>\leq\>\bm{\varepsilon}((T_{i})^{n_{i}})\>=\>\frac{\bm{\varepsilon}(T_{i})}{n_{i}}\>\leq\>\frac{10}{n_{i}\,rank(T_{i})},$
so that $n_{i}\,rank(T_{i})\leq 10\varepsilon^{-1}$ for each $i$. ∎
Fix $\varepsilon>0$. After Theorem A, it is natural to ask whether the fact that $G/Sol(G)$ is the direct product of boundedly many simple groups of Lie type of bounded rank, has some effect on the quasirandomness of the full $G$. In general, the answer is no, as it can happen that $G$ is perfect but its quasirandom degree vanishes when $|G|$ grows. It is enough to consider the wreath product $(C_{n})^{5}\rtimes\mathrm{Alt}(5)$ where $n$ is large: here $G/Sol(G)\cong\mathrm{Alt}(5)$ is a perfect group of bounded cardinality, so trivially quasirandom, while the small representations induced from $\mathrm{Alt}(5)$ destroy the quasirandomness of $G$. The following example, where $|G/Sol(G)|$ is unbounded, is due to S. Eberhard.
Example 4.5.
Let $p\geq 5$ be a prime, and $d\geq 2$. We define $G_{d}$ as the group of the $2d\times 2d$ upper triangular matrices whose diagonal entries are identical matrices in $\mathrm{SL}_{2}(p)$, and the upper entries are arbitrary $2\times 2$ matrices over $\mathbb{F}_{p}$. For instance,
$G_{3}\>=\>\left(\begin{array}[]{ccc}g&*&*\\ 0&g&*\\ 0&0&g\\ \end{array}\right),$
where $g\in\mathrm{SL}_{2}(p)$, and *** is arbitrary. We have
$|G_{d}|=\frac{1}{2}(p^{3}-p)p^{2d(d-1)}\geq p^{d^{2}}.$
It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix. In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $\frac{1}{4}$-quasirandom. On the other hand,
$\mathbf{D}_{0}(G_{d})\leq\mathbf{D}_{0}(\mathrm{PSL}_{2}(p))\leq p.$
It follows that $\bm{\varepsilon}(G_{d})\leq d^{-2}$, which vanishes when $d$ grows.
Despite the previous examples, a lower bound for $\bm{\varepsilon}(G)$ as a function of $\bm{\varepsilon}(G/Sol(G))$ exists when $|G/Sol(G)|$ is relatively large.
Lemma 4.6.
Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group. Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $\frac{\log|G/Sol(G)|}{\log|G|}\geq\delta$. If $|G|$ is sufficiently large, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
Proof.
By [8, Lemma 4.10], we have
$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
where $c$ is an absolute constant. Looking back at [18], we check that $c=10^{-10}$ works. It follows that
[…]
and passing to the logarithms,
$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/Sol(G))}{2}\cdot\frac{\log|G/Sol(G)|}{\log|G|}-\frac{10\log 10}{\log|G|}.$
The proof follows. ∎
4.2. Proof of Theorem B
We first prove the shallow version of Theorem B. This is a nice application of (2.1).
Proof of Corollary D.
From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree exceeds $\frac{1}{3}$. We claim that no other group has this property. Looking for a contradiction, suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$. Observe that $G$ is not simple, and let $N\lhd G$ be a nontrivial proper normal subgroup. We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$. In the former case, from (2.1), we have
[…]
So $N=1$, a contradiction. Now suppose $G/N\cong O^{\prime}N$. From (2.1),
[…]
so that $N$ is the cyclic group of order $2$. Since the only nonidentity element of $N$ has to be fixed by conjugation, $N$ is central in $G$. In particular, $G$ is a perfect central extension of $O^{\prime}N$ by $N$. On the other hand, the Schur multiplier of $O^{\prime}N$ has cardinality $3$, and such an extension does not exist. ∎
The following observation is not really needed in the proof of Theorem B, but it makes the structure of $G$ more transparent.
Lemma 4.7.
Suppose that $\bm{\varepsilon}(G)\geq\frac{1}{5}$, or that $\bm{\varepsilon}(G)\geq\frac{1}{6}$ and $|G|\geq 2\cdot 10^{24}$. Then $G$ has a unique maximal normal subgroup.
Proof.
From Corollary D, for every finite group $G$ we have
$\bm{\varepsilon}(G)\leq\max\left\{\bm{\varepsilon}(J_{1}),\bm{\varepsilon}(O^{\prime}N)\right\}=\bm{\varepsilon}(O^{\prime}N)<\frac{2}{5},$
while $\bm{\varepsilon}(G)<\frac{1}{3}$ if $G$ is not isomorphic to $J_{1}$ or $O^{\prime}N$. Suppose that $M_{1}$ and $M_{2}$ are two distinct maximal normal subgroups of the group $G$. Since $M_{1}M_{2}/M_{1}$ is a normal subgroup of the simple group $G/M_{1}$, we obtain that $G=M_{1}M_{2}$. So $M_{1}/(M_{1}\cap M_{2})$ and $M_{2}/(M_{1}\cap M_{2})$ are normal subgroups of $G/(M_{1}\cap M_{2})$ with trivial intersection. It follows that
$G/(M_{1}\cap M_{2})=M_{1}/(M_{1}\cap M_{2})\times M_{2}/(M_{1}\cap M_{2})\cong G/M_{2}\times G/M_{1}.$
Lemma 2.8 yields
$\bm{\varepsilon}(G)\leq\bm{\varepsilon}((G/M_{2})\times(G/M_{1}))<\frac{1}{5},$
and $\bm{\varepsilon}(G)<\frac{1}{6}$ if neither $J_{1}$ or $O^{\prime}N$ appears as a quotient of $G$. If one among $J_{1}$ or $O^{\prime}N$ appears as a quotient, then
[…]
because $\mathbf{D}_{0}(J_{1})<\mathbf{D}_{0}(O^{\prime}N)$. ∎
We are ready to prove Theorem B. This proof has a number of cases, the most difficult, which we treat in the end, being the following. We have an abelian minimal normal subgroup $N$ of $G$ such that $G/N$ is quasisimple, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of the simple quotient of $G/N$, say $L$. The key idea is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations. This gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We start with the details.
Remark 4.8.
The veracity of Theorem B has been checked with a computer for all the finite groups of order up to $2\cdot 10^{6}$, using the library PerfectGroups in the GAP System [20].
Let $G$ be a counterexample to Theorem B whose order is minimal. In particular, $\bm{\varepsilon}(G)\geq\frac{1}{5}$ and $G$ is not quasisimple. By Lemma 4.7, $G$ has a unique maximal normal subgroup and cannot be written as a direct product of nontrivial groups. Therefore, using also Lemma 2.12, we find a chain
$1\>\lhd\>N\>\lhd\>M\>\lhd\>G$
where $N$ is the unique minimal normal subgroup of $G$, and $M$ is the unique maximal normal subgroup of $G$. It follows that $G/M$ is isomorphic to a simple group $L$ satisfying $\bm{\varepsilon}(L)>\frac{1}{5}$, and that $N\cong T^{n}$ is the direct product of isomorphic simple groups. We distinguish two cases: either $G/N$ is quasisimple or not. Let us start from the latter. $G/N$ is not quasisimple. We have $\bm{\varepsilon}(G/N)>\frac{1}{5}$, and, by the minimality of $G$, we are in one of the two cases:
  1. $(a)$
    $G/N\cong(\mathbb{F}_{3})^{3}\rtimes\mathrm{SL}_{3}(3)$;
  2. $(b)$
    $G/N$ is a perfect central extension of $(\mathbb{F}_{q})^{4}\rtimes\mathrm{Sp}_{4}(q)$, $q=2^{f}$, $f\geq 6$.
If we are in case $(a)$, using (2.1) we obtain
[…]
Hence $N$ is trivial, a contradiction.
If we are in case $(b)$, we remark that $G/M\cong\mathrm{Sp}_{4}(q)$ and $\bm{\varepsilon}(G/N)<\frac{3}{14}$. Let $N\lhd N_{0}\lhd M$ such that $G/N_{0}\cong(\mathbb{F}_{q})^{4}\rtimes\mathrm{Sp}_{4}(q)$, $N_{0}/N\subseteq\mathbf{Z}(G/N)$. First suppose that $N$ is abelian. If $N\subseteq\mathbf{Z}(G)$, then $N_{0}\subseteq\mathbf{Z}(G)$ by Lemma 2.11. Thus $G$ is a perfect central extension of $(\mathbb{F}_{q})^{4}\rtimes\mathrm{Sp}_{4}(q)$, against the assumptions. If $N\nsubseteq\mathbf{Z}(G)$, then by Lemma 2.7,
$\frac{3}{14}>\bm{\varepsilon}(G/N)>\frac{\bm{\varepsilon}(G)}{1-\bm{\varepsilon}(G)}\geq\frac{1}{4}.$
This is impossible. It remains the case where $N$ is nonabelian, say $N=T^{n}$ where $T$ is a nonabelian finite simple group. We argue in a similar way as in Theorem A: we don’t have $C=\mathbf{C}_{G}(N)\subseteq N$ as before, but the same argument works because $C\subseteq M$. We have
(4.2) $|G|\leq\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
Considering the embedding of $G/C$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$. Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (4.2). $G/N$ is quasisimple. Let $C=\mathbf{C}_{G}(N)$. We distinguish three cases. Suppose that $C=G$. Here $N$ is central and $G$ is quasisimple by Lemma 2.10, a contradiction. Suppose that $C=1$. Here $N=T^{n}$ is the product of nonabelian finite simple groups, and $G$ is embedded into $\mathrm{Aut}(T^{n})$. Let $G/M=L$. Arguing as in the proof of Theorem A, we obtain
$|G|\leq\mathbf{D}_{0}(L)^{5},\hskip 14.22636pt\mbox{and}\hskip 14.22636pt|N|\geq 60^{\mathbf{D}_{0}(L)}.$
This is impossible for any value of $\mathbf{D}_{0}(L)$. Suppose that $N\subseteq C\subseteq M$. This case will take up the rest of the proof. First, Lemma 2.7 provides
$\bm{\varepsilon}(L)\geq\bm{\varepsilon}(G/N)>\frac{1}{4}.$
Moreover, $N$ is elementary abelian, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple). As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (2.1). We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one of them appears as $L$, then
[…]
and we can easily work wit a computer (see Remark 4.8). One last time, we divide the discussion into three cases:
  1. $(i)$
    $L$ is a simple group of Lie type defined over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in cross-characteristic;
  2. $(ii)$
    $L$ is a simple group of Lie type defined over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in natural characteristic;
  3. $(iii)$
    $L$ is a sporadic simple group.
Let $L$ be as described in $(i)$. We remark that $|L|>\mathbf{D}_{0}(L)^{3}$ by Corollary D. Hence, by (2.1),
(4.3) $|N|\leq|M|\leq\frac{\mathbf{D}_{0}(L)^{5}}{|L|}<\mathbf{D}_{0}(L)^{2}.$
From Remark 2.14, the definition of $\mathbf{R}_{r}(L)$ with $r\neq p$ a prime, and (4.3),
$2^{\mathbf{D}_{0}(L)-1}\leq 2^{\mathbf{R}_{r}(L)}\leq|N|<\mathbf{D}_{0}(L)^{2}.$
The above inequality is possible only if $\mathbf{D}_{0}(L)\leq 6$, which gives
[…]
This case can be handled computationally (see Remark 4.8). Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (2.1), we find
$\begin{split}|N|&\leq\frac{\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}}{|\mathrm{Sp}_{4}(q)|}\\ &\leq\frac{(q^{2}-1)^{5}(q^{2}-q)^{5}}{2^{5}(q+1)^{5}}\cdot\frac{1}{q^{4}(q^{4}-1)(q^{2}-1)}\\ &<q^{5}.\end{split}$
On the other hand, by Lemma 2.13 and Table 1,
$|N|\geq q^{4}.$
So $|N|=q^{4}$, and we are in the case of Theorem Bb.
The argument to deal with the other cases is the same, using (2.1), Lemma 2.13 and Table 1. When $L\cong\mathrm{PSL}_{2}(q)$,
$|N|\leq\frac{\mathbf{D}_{0}(\mathrm{PSL}_{2}(q))^{5}}{|\mathrm{PSL}_{2}(q)|}\leq\frac{(q-1)^{5}}{q(q^{2}-1)}<q^{2}.$
On the other hand,
$|N|\geq q^{2},$
which provides a contradiction. If $L\cong\mathrm{PSL}_{3}(q)$, $q\geq 4$, we obtain
$q^{3}\leq|N|\leq\frac{q^{5}(q+1)^{5}}{q^{3}(q^{3}-1)(q^{2}-1)(q-1)}<q^{3}.$
[…]
$q^{7}\leq|N|\leq\frac{(q+1)^{5}(q^{2}-q+1)^{5}}{q^{6}(q^{6}-1)(q^{2}-1)}<q^{7}.$
For ${}^{2}B_{2}(q^{2})$,
$q^{8}\leq|N|\leq\frac{q^{5}(q^{2}-1)^{5}}{2^{5/2}q^{4}(q^{4}-1)(q^{2}-1)}<q^{5}.$
Finally, for ${}^{2}G_{2}(q^{2})$,
$q^{14}\leq|N|\leq\frac{(q^{4}-q^{2}+1)^{5}}{q^{6}(q^{6}-1)(q^{2}-1)}<q^{7}.$
In all these four cases, the chain of inequalities is impossible. Let $L$ be as described in $(iii)$. Since $\bm{\varepsilon}(L)>\frac{1}{4}$, the possibilities for $L$ are
$M_{11},\quad J_{1},\quad J_{3},\quad^{2}F_{2}(4)^{\prime},\quad O^{\prime}N.$
Observe that $M_{11}$, $J_{1}$ and ${}^{2}F_{2}(4)^{\prime}$ admit no central perfect extensions, while the perfect central extensions of $J_{3}$ and $O^{\prime}N$ are not $\frac{1}{5}$-quasirandom. Therefore $N=M$ is the unique normal subgroup of $G$, and $N$ can be thought as the vector space of a nontrivial modular representation of $L$. For each sporadic group, we use [12] to obtain a lower bound on $|N|$, and we compare the result with (2.1). For $L\cong M_{11}$, we have
$3^{5}\leq|N|\leq\frac{10^{5}}{|M_{11}|}\leq 13;$
For $L\cong J_{1}$,
[…]
For $L\cong J_{3}$,
$3^{18}\leq|N|\leq\frac{85^{5}}{|J_{3}|}\leq 89;$
For $L\cong\>^{2}F_{2}(4)^{\prime}$,
$2^{26}\leq|N|\leq\frac{104^{5}}{|^{2}F_{2}(4)^{\prime}|}\leq 678;$
Finally, for $L\cong O^{\prime}N$,
[…]
These contradictions conclude the proof of Theorem B.
Appendix: Quasirandom degree of finite simple groups
We already discussed the alternating groups in Remark 4.4. In particular, we observed that they are the less quasirandom among the nonabelian finite simple groups.
In the following tables, we collect the values $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$ when $L$ is a finite simple group of Lie type or a sporadic simple group. While gathering these data, we stumbled into a little gap in the literature: as far as we know, the minimal degree of the complex irreducible representations for classical simple groups have never been computed and recorded in a single place. Let $L$ be a classical group of Lie type. In [28], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$ (that is, a complex irreducible representation of $\widetilde{L}$). In Table 2, we record $\mathbf{D}_{0}(L)$ for the classical groups of Lie type.
These data were virtually known, in the sense that all the complex representation of minimal dimension were already described as projective representations whose restrictions on the center are trivial. First, we notice that
$\mathrm{PSL}_{3}(2),\>\mathrm{PSL}_{3}(4),\>\mathrm{PSL}_{4}(2),\>\mathrm{PSL}_{4}(3),\>\mathrm{PSU}_{4}(2),\>\mathrm{PSU}_{4}(4),\>\Omega_{7}(3),\>\Omega^{+}_{8}(2)$
can be dealt with a calculator, while the character table of $\mathrm{PSL}_{2}(q)$ is well known. Now, note that unipotent representations of quasisimple groups are trivial on their center. Hence, using [28, Theorems 3.1, 4.1, 5.2, 7.6] and [17, Theorem 2.5], we obtain that for $\mathrm{PSL}_{d}(q)$, $\mathrm{PSU}_{d}(q)$ with $d$ odd, $\mathrm{Sp}_{2d}(q)$ with $q$ even, and $\mathrm{P}\Omega_{2d}^{\pm}(q)$, $\mathbf{D}_{0}(L)$ coincide with the minimal dimension of a projective unipotent representation. We are left to deal with $\mathrm{PSU}_{d}(q)$ with even rank, $\mathrm{PSp}_{2d}(q)$ in odd characteristic, and $\Omega_{2d+1}(q)$. We start with $\mathrm{PSU}_{d}(q)$ with $d$ even. By [28, Theorem 4.1], the second minimal dimension of a complex representation $\rho$ of $\mathrm{SU}_{d}(q)$ is
$\dim(\rho)=\mathbf{D}_{0}(\mathrm{SU}_{d}(q))+1=\frac{q^{d}+q}{q+1}.$
Further, such a representation is unipotent. Hence, $\mathbf{D}_{0}(\mathrm{PSU}_{d}(q))=\dim(\rho)$. Let us tackle $\mathrm{PSp}_{2d}(q)$ with $q$ odd. In view of [28, Theorem 5.2], the unitary representations $\rho_{+}$ and $\rho_{-}$ with smallest dimension of $\mathrm{Sp}_{2d}(q)$ are such that
$\dim(\rho_{-})=\frac{q^{d}-1}{2},\quad\hbox{and}\quad\dim(\rho_{+})=\frac{q^{d}+1}{2}.$
Observe that, as $\dim(\rho_{+})=\dim(\rho_{-})+1$, one of these representations has even degree, while the other has odd degree. The latter is trivial on the center, therefore it projects to an irreducible representation of $\mathrm{PSp}_{2d}(q)$ of minimal degree. Finally, we must consider $\Omega_{2d+1}(q)$. From [17, Theorem 2.2], the two unitary representations of $\mathrm{Spin}_{2d+1}(q)$ with minimal degrees are such that
$\dim(\rho_{1})=\frac{q^{2d}-1}{q^{2}-1},\quad\hbox{and}\quad\dim(\rho_{2})=\frac{(q^{d}-1)(q^{d}-q)}{2(q+1)}.$
Observe that $\rho_{2}$ is unipotent, while $\rho_{1}$ is not. For $q=3$, $\dim(\rho_{1})\leq\dim(\rho_{2})$, so we find that $\mathbf{D}_{0}(\Omega_{2d+1}(3))=\dim(\rho_{2})$. For all the other powers of odd primes, the inequality is reversed. As before, if $\dim(\rho_{1})$ is odd, $\rho_{1}$ is trivial on the center, thus $\mathbf{D}_{0}(\Omega_{2d+1}(q))=\dim(\rho_{1})$. Otherwise, we have $\mathbf{D}_{0}(\Omega_{2d+1}(q))=\dim(\rho_{2})$.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\mathbf{D}_{0}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d=2$, $q$ even $q-1$
[…]
$d\geq 3$, $(d,q)\neq(3,2),(4,2),(4,3)$ $\dfrac{q^{d}-1}{q-1}$
$d=3$, $q=2$ $3$
$d=4$, $q=2$ $7$
$d=4$, $q=3$ $26$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ odd $\dfrac{q^{d}-q}{q+1}$
$d\geq 4$ even, $(d,q)\neq(4,2)$ $\dfrac{q^{d}+q}{q+1}$
$d=4$, $q=2$ $5$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{2d}(q)$ $d\geq 2$, $q$ even $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q\neq 3$ and $12(q^{2}-1)\nmid q^{2d}-$ 1 $\dfrac{q^{2d}-1}{q^{2}-1}$
$d\geq 3$, $q=3$ or $12(q^{2}-1)\mid q^{2d}-$ 1 $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{-}_{2d}(q)$ $d\geq 4$ $\dfrac{(q^{d}+1)(q^{d-1}-q)}{q^{2}-1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{+}_{2d}(q)$ $d\geq 4$, $q\neq 2,3$ $\dfrac{(q^{d}-1)(q^{d-1}+q)}{q^{2}-1}$
$d=4$, $q=2$ $28$
$d\geq 5$, $q=2$ or $d\geq 4$, $q=3$ $\dfrac{(q^{d}-1)(q^{d-1}-1)}{q^{2}-1}$
Table 2. Minimal degree of a representation of simple classical groups.
In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups. These informations are collected from [7, 13, 14, 15, 17, 25, 27, 28].
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments Asymptotic for $|L|$ Asymptotic for $\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $q^{d^{2}-1}$ $q^{d-1}$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $q^{d^{2}-1}$ $q^{d-1}$ $\dfrac{1}{d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{2d}(q)$ $d\geq 2$, $q$ even $q^{2d^{2}+d}$ $q^{2d-1}$ $\dfrac{2d-1}{d(2d+1)}$
$d\geq 2$, $q$ odd $q^{2d^{2}+d}$ $q^{d}$ $\dfrac{1}{2d+1}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $q^{2d^{2}+d}$ $q^{2d-2}$ $\dfrac{2(d-1)}{d(2d+1)}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $q^{2d^{2}-d}$ $q^{2d-3}$ $\dfrac{2d-3}{d(2d-1)}$
Table 3. Asymptotic quasirandom degree of simple classical groups of Lie type.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments Asymptotic for $|L|$ Asymptotic for $\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{6}(q)$ $q^{78}$ $q^{11}$ $\dfrac{11}{78}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{7}(q)$ $q^{133}$ $q^{17}$ $\dfrac{17}{133}$
$\dfrac{c1^{1}}{1_{1_{1}}}$ $E_{8}(q)$ $q^{248}$ $q^{29}$ $\dfrac{29}{248}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ even $q^{52}$ $q^{11}$ $\dfrac{11}{52}$
$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $q^{52}$ $q^{8}$ $\dfrac{2}{13}$
[…]
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}E_{6}(q^{2})$ $q^{78}$ $q^{11}$ $\dfrac{11}{78}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{3}D_{4}(q^{3})$ $q^{28}$ $q^{5}$ $\dfrac{5}{28}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}F_{4}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $q^{52}$ $q^{11}$ $\dfrac{11}{52}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $q^{10}$ $q^{3}$ $\dfrac{3}{10}$
$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $q^{14}$ $q^{4}$ $\dfrac{2}{7}$
Table 4. Asymptotic quasirandom degree of simple exceptional groups of Lie type.
$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ $|L|$ $\mathbf{D}_{0}(L)$ $\bm{\varepsilon}(L)$ $L$ $|L|$ $\mathbf{D}_{0}(L)$ $\bm{\varepsilon}(L)$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{11}$ 7 92079207\,9207 920 $10$ $\approx 0.2565$ $Fi_{24}$ 125 520 $125\,520\,570\cdot 10^{16}$ 8 67186718\,6718 671 $\approx 0.1635$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{12}$ 95 0409504095\,04095 040 $11$ $\approx 0.2093$ $HS$ 44 352 0004435200044\,352\,00044 352 000 $22$ $\approx 0.1756$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{22}$ 443 520443520443\,520443 520 $21$ $\approx 0.2342$ $McL$ 898 128 000898128000898\,128\,000898 128 000 $22$ $\approx 0.1500$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{23}$ 10 200 9601020096010\,200\,96010 200 960 $22$ $\approx 0.1916$ $He$ 4 030 387 20040303872004\,030\,387\,2004 030 387 200 $51$ $\approx 0.1778$
$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{24}$ 244 823 040244823040244\,823\,040244 823 040 $23$ $\approx 0.1624$ $Ru$ 145 926 $145\,926\,144\cdot 10^{3}$ $378$ $\approx 0.2309$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{1}$ 175 560175560175\,560175 560 $56$ $\approx 0.3334$ $Suz$ 448 345 $448\,345\,497\cdot 10^{3}$ $143$ $\approx 0.1850$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{2}$ 604 800604800604\,800604 800 $14$ $\approx 0.1924$ $O^{\prime}N$ 460 815 $460\,815\,505\cdot 10^{3}$ 10 9441094410\,94410 944 $\approx 0.3464$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{3}$ 50 232 9605023296050\,232\,96050 232 960 $85$ $\approx 0.2506$ $HN$ 273 030 $273\,030\,912\cdot 10^{6}$ $133$ $\approx 0.1472$
$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{4}$ 867 755 $867\,755\,710\cdot 10^{11}$ 1 33313331\,3331 333 $\approx 0.1568$ $Ly$ 517 651 $517\,651\,790\cdot 10^{8}$ 2 48024802\,4802 480 $\approx 0.2031$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{3}$ 495 766 $495\,766\,656\cdot 10^{3}$ $23$ $\approx 0.1165$ $Th$ 907 459 $907\,459\,438\cdot 10^{8}$ $248$ $\approx 0.1413$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{2}$ 423 054 $423\,054\,213\cdot 10^{5}$ $23$ $\approx 0.1000$ $B$ 415 478 $415\,478\,148\cdot 10^{24}$ 4 37143714\,3714 371 $\approx 0.1083$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{1}$ 415 777 $415\,777\,680\cdot 10^{10}$ $276$ $\approx 0.1311$ $M$ 808 017 $808\,017\,424\cdot 10^{44}$ 196 883196883196\,883196 883 $\approx 0.0983$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{22}$ 645 617 $645\,617\,516\cdot 10^{5}$ $78$ $\approx 0.1371$ ${}^{2}F_{4}(2)^{\prime}$ 17 971 2001797120017\,971\,20017 971 200 $104$ $\approx 0.2781$
$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{23}$ 408 947 $408\,947\,047\cdot 10^{10}$ $782$ $\approx 0.1555$
Table 5. Quasirandom degree of the sporadic simple groups.

### v2

@@
 A more general discussion on quasirandomness is taken by Babai, Nikolov and Pyber [2], and Nikolov and Pyber [19].
-The celebrated Gower’s trick [18] says that, if $G$ is $\varepsilon$-quasirandom and $S\subseteq G$ is a large subset, namely $|S|\geq|G|^{1-\frac{\varepsilon}{3}}$, then $S^{3}=G$.
+The celebrated Gowers’ trick [19] says that, if $G$ is $\varepsilon$-quasirandom and $S\subseteq G$ is a large subset, namely $|S|\geq|G|^{1-\frac{\varepsilon}{3}}$, then $S^{3}=G$.
 This fact, as well as other natural invariants of $G$, can be used to characterize quasirandomness.
@@
 The present article was born from our need to answer the question: what, concretely, is a quasirandom group?
-In the first main result, we prove that large $\varepsilon$-quasirandom groups with a trivial solvable radical are direct products of nonabelian simple groups.
-We say a little more.
+In the first main result we describe the large $\varepsilon$-quasirandom groups, for any fixed $\varepsilon>0$.
+They are precisely the groups whose quotient by the solvable radical is a relatively large direct product of boundedly many finite simple groups of Lie type of bounded rank.
 Theorem A.
-Fix $\varepsilon>0$, and let $G$ be an $\varepsilon$-quasirandom group.
+Fix $\varepsilon>0$.
+If $G$ is an $\varepsilon$-quasirandom group, then $|G/Sol(G)|\geq|G|^{\varepsilon}$.
 If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/Sol(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
+Conversely, suppose that $G$ is perfect, $|G/Sol(G)|\geq|G|^{\varepsilon}$, and that $G/Sol(G)$ is $\varepsilon$-quasirandom.
+If $|G|$ is sufficiently large depending on $\varepsilon$, then $G$ is $\frac{\varepsilon^{2}}{3}$-quasirandom.
 For the second main result, we study the groups that are $\varepsilon$-quasirandom with a relatively large $\varepsilon$.
-It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the most quasirandom family of groups [24, 29], but not much is known.
+It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the “most quasirandom” family of groups [23, 28], but not much is known.
 Not only this is actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
@@
     $G$ is quasisimple.
-The irreducible representations of the quasisimple groups are known [27], thus it is an easy (although laborious) task to recover the full list of the $\frac{1}{5}$-quasirandom groups from Theorem B.
+The low dimensional representations of the quasisimple groups are known [26], thus it is an easy (although laborious) task to recover the full list of the $\frac{1}{5}$-quasirandom groups from Theorem B.
 An interesting consequence is the following:
@@
 It also makes the proof a little cleaner, because every $\frac{1}{5}$-quasirandom group has a unique maximal normal subgroup (Lemma 4.8).
-We point out that similar methods are useful to study quasirandomnes at other regimes different from the exponential, which could be of interest as well.
+We point out that similar methods are useful to study quasirandomnes at other regimes different from the exponential, which could be of interest as well (see also Remark 4.6).
 The shallow version of Theorem B is the following:
@@
 The philosophy behind the proofs is that quasirandom groups are pretty rigid, and lend themselves very little to nontrivial extensions, especially the noncentral ones.
-To make this idea really say something, we use various properties of the finite simple groups and their representations, both complex and modular, linear and projective.
+To make this idea really say something, we use various properties of the finite simple groups and their representations.
+Particularly for Theorem B, we need a detailed (i.e.
+quantitative) analysis of quasirandom degrees, modular and projective representations.
 The paper is structured as follows.
@@
 We prove the main theorems in Section 4.
-Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups: we collect these data in the Appendix.
+Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups, so we collect these data in the Appendix.
 2. Preliminaries
@@
 $\mathbf{D}_{0}(G)\>:=\>\min\left\{\dim(\pi)\mid\pi\in\mathrm{Irr}(G)\setminus\{1\}\right\}.$
-We start with the following, which appears as [26, Ex. 1.3.2].
+We start with the following, which appears as [25, Ex. 1.3.2].
 Lemma 2.1.
 Let $N\lhd G$ be a nontrivial proper normal subgroup. Then
-$\mathbf{D}_{0}(G/N)\,\geq\,\mathbf{D}_{0}(G)\,.$
+$\mathbf{D}_{0}(G/N)\>\geq\>\mathbf{D}_{0}(G).$
 If the previous inequality is strict, then
@@
 ∎
-We would like to mention some connections between $\mathbf{D}_{0}(G)$ and other important parameters associated to a finite group: the number of conjugacy classes $\mathbf{k}(G)$, the representation growth $\mathbf{r}_{n}(G)$, and the minimal index of a proper subgroup, namely $\bm{\iota}(G):=\min_{H<G}|G:H|$.
+We would like to mention some connections between $\mathbf{D}_{0}(G)$ and other important parameters associated to a finite group: the number of conjugacy classes $\mathbf{k}(G)$, the representation growth $\mathbf{r}_{n}(G)$, and the minimal index of a proper subgroup, i.e.
+$\bm{\iota}(G):=\min_{H<G}|G:H|$.
 Remark 2.4.
@@
 $|G|=\sum_{\pi\in\mathrm{Irr}(G)}\dim(\pi)^{2}\geq 1+(\mathbf{k}(G)-1)\mathbf{D}_{0}(G)^{2}.$
-It follows that, if $G$ is an $\varepsilon$-quasirandom group, then $\mathbf{k}(G)\leq|G|^{1-2\varepsilon}$.
-Moreover, the number $\mathbf{r}_{n}(G)$ of the inequivalent irreducible representations of $G$ of degree $n$ is polynomially bounded, i.e.
-$\mathbf{r}_{n}(G)\leq n^{\varepsilon^{-1}-2}$ for all $n$ (see [22, Lemma 25]).
+Let $G$ be an $\varepsilon$-quasirandom group.
+It follows that $\varepsilon<\frac{1}{2}$, and $\mathbf{k}(G)\leq|G|^{1-2\varepsilon}$.
+Moreover, the number $\mathbf{r}_{n}(G)$ of the inequivalent irreducible representations of degree $n$ is polynomially bounded, namely $\mathbf{r}_{n}(G)\leq n^{\varepsilon^{-1}-2}$ for all $n$ (see [21, Lemma 25]).
 Remark 2.5.
@@
 Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$).
-In particular, $\mathbf{D}_{0}(G)<|G|^{0.5}$.
-As consequences of Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
+From Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
 2.2. Quasirandom degree
@@
 $\bm{\varepsilon}(G)\>:=\>\log_{|G|}\mathbf{D}_{0}(G).$
-By definition, $\mathbf{D}_{0}(G)=|G|^{\bm{\varepsilon}(G)}$, and $\bm{\varepsilon}(G)$ is the maximum $\varepsilon\in[0,1)$ such that $G$ is an $\varepsilon$-quasirandom group.
+By definition, $\mathbf{D}_{0}(G)=|G|^{\bm{\varepsilon}(G)}$, and $\bm{\varepsilon}(G)$ is the maximum $\varepsilon\in[0,\frac{1}{2})$ such that $G$ is an $\varepsilon$-quasirandom group.
 Observe that $\bm{\varepsilon}(G)>0$ if and only if $G$ is perfect.
@@
 Lemma 2.7.
-Let $N\lhd G$ be a nontrivial noncentral abelian normal subgroup. Then
+Let $N\lhd G$ be a noncentral abelian normal subgroup. Then
 $\bm{\varepsilon}(G/N)\>>\>\frac{\bm{\varepsilon}(G)}{1-\bm{\varepsilon}(G)}.$
@@
 A finite group is quasisimple if it is a perfect central extension of a nonabelian finite simple group, that is, $G$ is perfect and $G/\mathbf{Z}(G)$ is simple.
-It is a key property of perfect groups that, if $G$ is perfect, then there is a unique universal central perfect extension, which we denote by $\widetilde{G}$.
+It is a key property of perfect groups that, if $G$ is perfect, then there is a unique universal perfect central extension of $G$, which we denote by $\widetilde{G}$.
 This extension is obtained by choosing $N$ isomorphic to the Schur multiplier of $G$, that is, the second homology group $H^{2}(G,\mathbb{Z})$.
@@
 2.4. Linear and projective representations
-So far, we have only dealt with complex linear representations, but modular and projective representations will be a key tool in the proof of the main theorems.
+So far, we have only dealt with complex linear representations, but modular and projective representations will be a key tool in the proof of Theorem B.
 For a prime $r\geq 2$, we use $\mathbf{D}_{r}(G)$ to denote the minimal degree of a nontrivial irreducible modular representation of $G$ over a field of characteristic $r$.
@@
 The degree of the representation is the dimension of $V$.
-As for linear representations, we denote by $\mathbf{R}_{0}(G)$ the minimal degree of a nontrivial projective complex representation, and by $\mathbf{R}_{r}(G)$ the minimal degree of a nontrivial projective representation over a field of characteristic $r$.
+As for linear representations, we denote by $\mathbf{R}_{0}(G)$ the minimal degree of a nontrivial complex projective representation, and by $\mathbf{R}_{r}(G)$ the minimal degree of a nontrivial projective representation over a field of characteristic $r$.
 We remark that, if $G$ is finite and perfect, and $\widetilde{G}$ is its universal perfect central extension, then
@@
 In the remainder of this subsection, we focus on the projective representations of finite simple groups of Lie type.
-We recall that, if $G$ is a group of Lie type defined over a field of characteristic $p$, then a modular representation of $G$ is in natural characteristic when the characteristic of the underlying vector space is $p$, while it is in cross-characteristic otherwise.
+We recall that, if $G$ is a group of Lie type over a field of characteristic $p$, then a modular representation of $G$ over a vector space $V$ is in natural characteristic when the characteristic of $V$ is $p$, while it is in cross-characteristic otherwise.
 Lemma 2.13.
@@
 Proof.
-It follows from Propositions 5.3.1, 5.4.6, and Remark 5.4.7 of [13]. ∎
-For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom simple groups. The infinite families of $\frac{1}{4}$-quasirandom groups are
+It follows immediately from the definition. ∎
+For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom finite simple groups. The infinite families of $\frac{1}{4}$-quasirandom simple groups are
 $\mathrm{PSL}_{d}(q)\hbox{ with }d\in\{2,3\},\quad\mathrm{Sp}_{4}(q)\hbox{ with }q\hbox{ even},$
 $G_{2}(q)\hbox{ with }q=3(\mathrm{mod}\,6),\quad^{2}B_{2}(q^{2}),\quad^{2}G_{2}(q^{2})\,.$
-The relevant informations can be recovered from [14, 25] for classical groups, and from [15, Sec. 4] for exceptional groups. We collect these data in Table 1.
+The relevant informations can be recovered from [15, 24] for classical groups, and from [17, Sec. 4] for exceptional groups (see also the Appendix). We collect these data in Table 1.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\mathbf{D}_{0}(L)$ $\mathbf{R}_{0}(L)$ $\mathbf{R}_{r}(L)$, $r\neq p$ $\mathbf{R}_{p}(L)$
@@
 $\mathbf{R}_{r}(L)\>\geq\>\mathbf{D}_{0}(L)-1.$
-Moreover, we point out that the comparison between $\mathbf{R}_{0}(L)$ and $\mathbf{R}_{r}(L)$ can be explicitly found, for all simple groups of Lie type, in [27, Sec. 3].
+Moreover, we point out that the comparison between $\mathbf{R}_{0}(L)$ and $\mathbf{R}_{r}(L)$ ($r\neq p$) can be explicitly found, for all simple groups of Lie type, in [26, Sec. 3].
 3. Quasirandom groups of affine type
@@
 The key tool is Lemma 3.2 below.
-Before giving the results, we revise some notions from the representation theory of abelian groups.
+Before giving the results, we revise some notions from the representation theory of finite groups.
 Let $V$ be a finite vector space, and let $V^{*}$ be its dual space.
@@
 $\rho^{*}(g)\>:=\>\rho(g^{-1})^{*}$
-for every $g\in K$.
+for every $g\in K$. It is well known that the representations $\rho$ and $\rho^{*}$ are not isomorphic in general.
 Lemma 3.1.
@@
 The number of $g$-fixed points in $V$ is $q^{\dim(Ker(g-I))}$, while the number of $g$-fixed points in $V^{*}$ is $q^{\dim(Ker((g^{-1})^{T}-I))}$, where $g^{T}$ is the transpose of $g$.
-But $\dim(Ker(g-I))$ is the multiplicity of $1$ as an eigenvalue of $g$, so the two quantities above are equal.
-Since $g$ is arbitrary, the proof follows from Burnside’s lemma.
+Now $\dim(Ker(g-I))$ is the multiplicity of $1$ as an eigenvalue of $g$.
+We remark that the eigenvalues of $(g^{-1})^{T}$ are the eigenvalues of $g^{-1}$, which are the inverses of the eigenvalues of $g$.
+In particular, the two quantities above are equal.
+Since $g$ is arbitrary, the proof follows from Burnside’s lemma, i.e.
+the number of orbits is equal to the average number of fixed points.
 ∎
@@
 4.1. Proof of Theorem A
-The main idea behind the proof of Theorem A is the following.
+We immediately prove the second part of the theorem. We restate it as follows.
+Lemma 4.1.
+Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group.
+Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $|G/Sol(G)|\geq|G|^{\delta}$.
+If $|G|$ is sufficiently large with respect to $\varepsilon$ and $\delta$, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
+Proof.
+By [8, Lemma 4.10], we have
+$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
+where $c$ is an absolute constant. Looking back at [19], we check that $c=10^{-10}$ works. It follows that
+[…]
+and passing to the logarithms,
+$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/Sol(G))}{2}\cdot\frac{\log|G/Sol(G)|}{\log|G|}-\frac{10\log 10}{\log|G|}.$
+The proof follows. ∎
+The main idea behind the other direction of Theorem A is the following.
 Under the hypothesis that $Sol(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$.
@@
 In the following remark, we discuss the asymptotic behaviour of $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$, when $L$ is a finite simple group, and $|L|\rightarrow+\infty$.
+Of course, we do not care about the sporadic simple groups in this context.
 Remark 4.5.
@@
 Proof of Theorem A.
+After Lemma 4.1, we have only to prove the first part of the theorem.
 Fix $\varepsilon>0$.
-From Corollary 2.3(i), it is clear that if $|G|$ is large then $|G/Sol(G)|$ is large.
+From Corollary 2.3(i), it is clear that $|G/Sol(G)|\geq|G|^{\varepsilon}$.
 Moreover, $G/Sol(G)$ is again $\varepsilon$-quasirandom.
@@
 Otherwise, it is easy to see that $G/F$ has a simple quotient $L$ which is $\varepsilon$-quasirandom.
-Of course, $|G|\leq\mathbf{D}_{0}(L)^{1/\varepsilon}$.
-So, if $|L|$ is bounded, then $|G|$ is bounded.
-It follows that $L=L_{d}(q)$ is a simple group of Lie type, $d$ is $\varepsilon$-bounded, while $q$ is unbounded.
 By the definition of quasirandom degree, we have
-(4.1) $|G|\>\leq\>\mathbf{D}_{0}(L_{d}(q))^{1/\varepsilon}.$
+(4.1) $|G|\>\leq\>\mathbf{D}_{0}(L)^{1/\varepsilon}.$
+So, if $\mathbf{D}_{0}(L)$ is bounded, then $|G|$ is bounded.
 Let $n=\max_{i}n_{i}$, so that $|F|\geq 60^{n}$.
-From Remark 4.2 and Theorem 4.1, $L_{d}(q)$ has to appear as a section of $\mathrm{Alt}(n)$.
-From Lemma 4.3, $\mathrm{Sym}(n)$ has a subgroup isomorphic to $L_{d}(q)$.
-From Lemma 2.2 we obtain $\mathbf{D}_{0}(L_{d}(q))\leq n$, and so
+From Remark 4.3 and Theorem 4.2, $L$ has to appear as a section of $\mathrm{Alt}(n)$.
+From Lemma 4.4, $\mathrm{Sym}(n)$ has a subgroup isomorphic to $L$.
+From Lemma 2.2 we obtain $\mathbf{D}_{0}(L)\leq n$, and so
 […]
-When $\mathbf{D}_{0}(L_{d}(q))$ is sufficiently large, this gives a contradiction to (4.1).
-But $\mathbf{D}_{0}(L_{d}(q))\geq|G|^{\varepsilon}$, so we obtain a contradiction when $|G|$ is sufficiently large.
+When $\mathbf{D}_{0}(L)$ is sufficiently large, this gives a contradiction to (4.1).
 We have proven that $G/Sol(G)=F^{*}(G/Sol(G))$, so $G/Sol(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/Sol(G)=\prod_{i}(T_{i})^{n_{i}}$.
@@
 so that $n_{i}\,rank(T_{i})\leq 10\varepsilon^{-1}$ for each $i$. ∎
-Fix $\varepsilon>0$.
-After Theorem A, it is natural to ask whether the fact that $G/Sol(G)$ is the direct product of boundedly many simple groups of Lie type of bounded rank, has some effect on the quasirandomness of the full $G$.
-In general, the answer is no, as it can happen that $G$ is perfect but its quasirandom degree vanishes when $|G|$ grows.
-It is enough to consider the wreath product $(C_{n})^{5}\rtimes\mathrm{Alt}(5)$ where $n$ is large: here $G/Sol(G)\cong\mathrm{Alt}(5)$ is a perfect group of bounded cardinality, so trivially quasirandom, while the small representations induced from $\mathrm{Alt}(5)$ destroy the quasirandomness of $G$.
-The following example, where $|G/Sol(G)|$ is unbounded, is due to S.
+Remark 4.6.
+An attentive reader will have noticed that a bound of the type $\mathbf{D}_{0}(G)\geq c\log|G|$ is strong enough to give that $G/Sol(G)$ is the direct product of nonabelian finite simple groups.
+It is natural to ask whether the condition $|G/Sol(G)|\geq|G|^{\varepsilon}$ is actually superflous in the second part of Theorem A, as it may be implied by the quasirandomness of the sole $G/Sol(G)$.
+The answer is no, as it can even happen that $|G/Sol(G)|$ is bounded: it is enough to consider the perfect deleted permutation module $(C_{n})^{4}\rtimes\mathrm{Alt}(5)$ when $n$ is large.
+The following example, where $|G/Sol(G)|$ is unbounded but too small, is due to S.
 Eberhard.
@@
 $G_{3}\>=\>\left(\begin{array}[]{ccc}g&*&*\\ 0&g&*\\ 0&0&g\\ \end{array}\right),$
-where $g\in\mathrm{SL}_{2}(p)$, and *** is arbitrary. We have
+where $g\in\mathrm{SL}_{2}(p)$, and * is arbitrary. We have
 $|G_{d}|=\frac{1}{2}(p^{3}-p)p^{2d(d-1)}\geq p^{d^{2}}.$
 It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix.
-In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $\frac{1}{4}$-quasirandom.
-On the other hand,
-$\mathbf{D}_{0}(G_{d})\leq\mathbf{D}_{0}(\mathrm{PSL}_{2}(p))\leq p.$
-It follows that $\bm{\varepsilon}(G_{d})\leq d^{-2}$, which vanishes when $d$ grows.
-Despite the previous examples, a lower bound for $\bm{\varepsilon}(G)$ as a function of $\bm{\varepsilon}(G/Sol(G))$ exists when $|G/Sol(G)|$ is relatively large.
-Lemma 4.6.
-Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group.
-Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $\frac{\log|G/Sol(G)|}{\log|G|}\geq\delta$.
-If $|G|$ is sufficiently large, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
-Proof.
-By [8, Lemma 4.10], we have
-$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
-where $c$ is an absolute constant. Looking back at [18], we check that $c=10^{-10}$ works. It follows that
-[…]
-and passing to the logarithms,
-$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/Sol(G))}{2}\cdot\frac{\log|G/Sol(G)|}{\log|G|}-\frac{10\log 10}{\log|G|}.$
-The proof follows. ∎
+In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $0.34$-quasirandom for large $p$, but its cardinality is less than $p^{3}$.
 4.2. Proof of Theorem B
@@
 Proof of Corollary D.
-From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree exceeds $\frac{1}{3}$.
-We claim that no other group has this property.
-Looking for a contradiction, suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$.
-Observe that $G$ is not simple, and let $N\lhd G$ be a nontrivial proper normal subgroup.
+From the data collected in the Appendix, we note that, among the finite simple groups, $J_{1}$ and $O^{\prime}N$ are the only ones whose quasirandom degree is at least $\frac{1}{3}$.
+Suppose that $G$ is of minimal cardinality among the finite groups satisfying $\bm{\varepsilon}(G)\geq\frac{1}{3}$ that are not isomorphic to either $J_{1}$ or $O^{\prime}N$.
+Of course, there exists a nontrivial proper normal subgroup $N$ of $G$.
 We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$.
@@
 We have an abelian minimal normal subgroup $N$ of $G$ such that $G/N$ is quasisimple, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of the simple quotient of $G/N$, say $L$.
-The key idea is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations.
+The key point is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations.
 This gives a lower bound on $|N|$, which we put against the upper bound given by (2.1).
@@
 Let us start from the latter.
-$G/N$ is not quasisimple.
-We have $\bm{\varepsilon}(G/N)>\frac{1}{5}$, and, by the minimality of $G$, we are in one of the two cases:
+$G/N$ is not quasisimple. We have $\bm{\varepsilon}(G/N)>\frac{1}{5}$, and, by the minimality of $G$, we are in one of the two cases:
   1. $(a)$
@@
 It remains the case where $N$ is nonabelian, say $N=T^{n}$ where $T$ is a nonabelian finite simple group.
-We argue in a similar way as in Theorem A: we don’t have $C=\mathbf{C}_{G}(N)\subseteq N$ as before, but the same argument works because $C\subseteq M$.
+We argue in a similar way as in Theorem A: since $N$ is the unique minimal normal subgroup of $G$, we have $\mathbf{C}_{G}(N)=1$ and the same argument works.
 We have
-(4.2) $|G|\leq\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
-Considering the embedding of $G/C$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$.
+(4.2) $|G|\>\leq\>\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
+Considering the embedding of $G$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$.
 Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (4.2).
-$G/N$ is quasisimple.
-Let $C=\mathbf{C}_{G}(N)$.
-We distinguish three cases.
-Suppose that $C=G$.
-Here $N$ is central and $G$ is quasisimple by Lemma 2.10, a contradiction.
-Suppose that $C=1$.
-Here $N=T^{n}$ is the product of nonabelian finite simple groups, and $G$ is embedded into $\mathrm{Aut}(T^{n})$.
-Let $G/M=L$.
-Arguing as in the proof of Theorem A, we obtain
+$G/N$ is quasisimple. Let $C=\mathbf{C}_{G}(N)$. We distinguish three cases.
+Suppose that $C=G$. Here $N$ is central and $G$ is quasisimple by Lemma 2.10, a contradiction.
+Suppose that $C=1$. Here $N=T^{n}$ is the product of nonabelian finite simple groups, and $G$ is embedded into $\mathrm{Aut}(T^{n})$. Let $G/M=L$. Arguing again as in the proof of Theorem A, we obtain
 $|G|\leq\mathbf{D}_{0}(L)^{5},\hskip 14.22636pt\mbox{and}\hskip 14.22636pt|N|\geq 60^{\mathbf{D}_{0}(L)}.$
-This is impossible for any value of $\mathbf{D}_{0}(L)$. Suppose that $N\subseteq C\subseteq M$. This case will take up the rest of the proof. First, Lemma 2.7 provides
+This is impossible for any value of $\mathbf{D}_{0}(L)$.
+Suppose that $N\subseteq C\subseteq M$. This case will take up the rest of the proof. First, $N$ is an elementary abelian noncentral subgroup, and Lemma 2.7 provides
 $\bm{\varepsilon}(L)\geq\bm{\varepsilon}(G/N)>\frac{1}{4}.$
-Moreover, $N$ is elementary abelian, and the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple).
+Moreover, the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple).
 As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (2.1).
-We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one of them appears as $L$, then
+We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one among them appears as $L$, then
 […]
@@
   1. $(i)$
-    $L$ is a simple group of Lie type defined over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in cross-characteristic;
+    $L$ is a simple group of Lie type over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in cross-characteristic;
   2. $(ii)$
-    $L$ is a simple group of Lie type defined over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in natural characteristic;
+    $L$ is a simple group of Lie type over a field of characteristic $p$, and $G/N$ induces a projective modular representation of $L$ in natural characteristic;
   3. $(iii)$
@@
 Let $L$ be as described in $(i)$. We remark that $|L|>\mathbf{D}_{0}(L)^{3}$ by Corollary D. Hence, by (2.1),
-(4.3) $|N|\leq|M|\leq\frac{\mathbf{D}_{0}(L)^{5}}{|L|}<\mathbf{D}_{0}(L)^{2}.$
-From Remark 2.14, the definition of $\mathbf{R}_{r}(L)$ with $r\neq p$ a prime, and (4.3),
+$|N|\leq|M|\leq\frac{\mathbf{D}_{0}(L)^{5}}{|L|}<\mathbf{D}_{0}(L)^{2}.$
+From Remark 2.14, and the definition of $\mathbf{R}_{r}(L)$ with $r\neq p$ a prime, we obtain
 $2^{\mathbf{D}_{0}(L)-1}\leq 2^{\mathbf{R}_{r}(L)}\leq|N|<\mathbf{D}_{0}(L)^{2}.$
@@
 […]
-This case can be handled computationally (see Remark 4.8).
-Let $L$ be as described in $(ii)$.
-We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$.
-Since the Schur multiplier is trivial, we have $N=M$.
-By (2.1), we find
+This case can be handled computationally.
+Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (2.1), we find
 $\begin{split}|N|&\leq\frac{\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}}{|\mathrm{Sp}_{4}(q)|}\\ &\leq\frac{(q^{2}-1)^{5}(q^{2}-q)^{5}}{2^{5}(q+1)^{5}}\cdot\frac{1}{q^{4}(q^{4}-1)(q^{2}-1)}\\ &<q^{5}.\end{split}$
@@
 $|N|\geq q^{4}.$
-So $|N|=q^{4}$, and we are in the case of Theorem Bb.
+So $|N|=q^{4}$, and we are in the situation of Theorem Bb.
 The argument to deal with the other cases is the same, using (2.1), Lemma 2.13 and Table 1. When $L\cong\mathrm{PSL}_{2}(q)$,
@@
 $q^{14}\leq|N|\leq\frac{(q^{4}-q^{2}+1)^{5}}{q^{6}(q^{6}-1)(q^{2}-1)}<q^{7}.$
-In all these four cases, the chain of inequalities is impossible. Let $L$ be as described in $(iii)$. Since $\bm{\varepsilon}(L)>\frac{1}{4}$, the possibilities for $L$ are
+In all these four cases, the chain of inequalities is impossible.
+Let $L$ be as described in $(iii)$. Since $\bm{\varepsilon}(L)>\frac{1}{4}$, the possibilities for $L$ are
 $M_{11},\quad J_{1},\quad J_{3},\quad^{2}F_{2}(4)^{\prime},\quad O^{\prime}N.$
@@
 In the following tables, we collect the values $\mathbf{D}_{0}(L)$ and $\bm{\varepsilon}(L)$ when $L$ is a finite simple group of Lie type or a sporadic simple group.
-While gathering these data, we stumbled into a little gap in the literature: as far as we know, the minimal degree of the complex irreducible representations for classical simple groups have never been computed and recorded in a single place.
-Let $L$ be a classical group of Lie type.
-In [28], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$ (that is, a complex irreducible representation of $\widetilde{L}$).
-In Table 2, we record $\mathbf{D}_{0}(L)$ for the classical groups of Lie type.
-These data were virtually known, in the sense that all the complex representation of minimal dimension were already described as projective representations whose restrictions on the center are trivial.
-First, we notice that
-$\mathrm{PSL}_{3}(2),\>\mathrm{PSL}_{3}(4),\>\mathrm{PSL}_{4}(2),\>\mathrm{PSL}_{4}(3),\>\mathrm{PSU}_{4}(2),\>\mathrm{PSU}_{4}(4),\>\Omega_{7}(3),\>\Omega^{+}_{8}(2)$
-can be dealt with a calculator, while the character table of $\mathrm{PSL}_{2}(q)$ is well known.
-Now, note that unipotent representations of quasisimple groups are trivial on their center.
-Hence, using [28, Theorems 3.1, 4.1, 5.2, 7.6] and [17, Theorem 2.5], we obtain that for $\mathrm{PSL}_{d}(q)$, $\mathrm{PSU}_{d}(q)$ with $d$ odd, $\mathrm{Sp}_{2d}(q)$ with $q$ even, and $\mathrm{P}\Omega_{2d}^{\pm}(q)$, $\mathbf{D}_{0}(L)$ coincide with the minimal dimension of a projective unipotent representation.
-We are left to deal with $\mathrm{PSU}_{d}(q)$ with even rank, $\mathrm{PSp}_{2d}(q)$ in odd characteristic, and $\Omega_{2d+1}(q)$.
-We start with $\mathrm{PSU}_{d}(q)$ with $d$ even.
-By [28, Theorem 4.1], the second minimal dimension of a complex representation $\rho$ of $\mathrm{SU}_{d}(q)$ is
-$\dim(\rho)=\mathbf{D}_{0}(\mathrm{SU}_{d}(q))+1=\frac{q^{d}+q}{q+1}.$
-Further, such a representation is unipotent.
-Hence, $\mathbf{D}_{0}(\mathrm{PSU}_{d}(q))=\dim(\rho)$.
-Let us tackle $\mathrm{PSp}_{2d}(q)$ with $q$ odd.
-In view of [28, Theorem 5.2], the unitary representations $\rho_{+}$ and $\rho_{-}$ with smallest dimension of $\mathrm{Sp}_{2d}(q)$ are such that
-$\dim(\rho_{-})=\frac{q^{d}-1}{2},\quad\hbox{and}\quad\dim(\rho_{+})=\frac{q^{d}+1}{2}.$
-Observe that, as $\dim(\rho_{+})=\dim(\rho_{-})+1$, one of these representations has even degree, while the other has odd degree.
-The latter is trivial on the center, therefore it projects to an irreducible representation of $\mathrm{PSp}_{2d}(q)$ of minimal degree.
-Finally, we must consider $\Omega_{2d+1}(q)$.
-From [17, Theorem 2.2], the two unitary representations of $\mathrm{Spin}_{2d+1}(q)$ with minimal degrees are such that
-$\dim(\rho_{1})=\frac{q^{2d}-1}{q^{2}-1},\quad\hbox{and}\quad\dim(\rho_{2})=\frac{(q^{d}-1)(q^{d}-q)}{2(q+1)}.$
-Observe that $\rho_{2}$ is unipotent, while $\rho_{1}$ is not.
-For $q=3$, $\dim(\rho_{1})\leq\dim(\rho_{2})$, so we find that $\mathbf{D}_{0}(\Omega_{2d+1}(3))=\dim(\rho_{2})$.
-For all the other powers of odd primes, the inequality is reversed.
-As before, if $\dim(\rho_{1})$ is odd, $\rho_{1}$ is trivial on the center, thus $\mathbf{D}_{0}(\Omega_{2d+1}(q))=\dim(\rho_{1})$.
-Otherwise, we have $\mathbf{D}_{0}(\Omega_{2d+1}(q))=\dim(\rho_{2})$.
+For a group of Lie type $L_{d}(q)$, we fix $d$ and make $q$ go to infinity.
+The case where $L$ is of exceptional type is discussed in detail by Lübeck [17].
+So let $L$ be a classical group of Lie type.
+In [27], Tiep and Zalesskii computed the minimal dimension of a projective complex irreducible representation of $L$, which we denote by $\mathbf{R}_{0}(L)$.
+See also [18].
+In Table 2, we use these informations to record $\mathbf{D}_{0}(L)$ for the simple classical groups.
+We remark that $\mathbf{R}_{0}(L)<\mathbf{D}_{0}(L)$ holds in some cases, for example for some symplectic groups in odd characteristic.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\mathbf{D}_{0}(L)$
@@
 […]
-$d\geq 3$, $(d,q)\neq(3,2),(4,2),(4,3)$ $\dfrac{q^{d}-1}{q-1}$
+$d\geq 3$, $(d,q)\neq(3,2),(4,2),(4,3)$ $\dfrac{q^{d}-q}{q-1}$
 $d=3$, $q=2$ $3$
@@
 $\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ odd $\dfrac{q^{d}-q}{q+1}$
-$d\geq 4$ even, $(d,q)\neq(4,2)$ $\dfrac{q^{d}+q}{q+1}$
+$d\geq 4$ even, $(d,q)\neq(4,2)$ $\dfrac{q^{d}-1}{q+1}\leq\mathbf{D}_{0}(L)\leq\dfrac{q^{d}+q}{q+1}$
 $d=4$, $q=2$ $5$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{2d}(q)$ $d\geq 2$, $q$ even $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
 […]
@@
 $d\geq 5$, $q=2$ or $d\geq 4$, $q=3$ $\dfrac{(q^{d}-1)(q^{d-1}-1)}{q^{2}-1}$
-Table 2. Minimal degree of a representation of simple classical groups.
-In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups.
-These informations are collected from [7, 13, 14, 15, 17, 25, 27, 28].
-$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments Asymptotic for $|L|$ Asymptotic for $\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $q^{d^{2}-1}$ $q^{d-1}$ $\dfrac{1}{d+1}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $q^{d^{2}-1}$ $q^{d-1}$ $\dfrac{1}{d+1}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{2d}(q)$ $d\geq 2$, $q$ even $q^{2d^{2}+d}$ $q^{2d-1}$ $\dfrac{2d-1}{d(2d+1)}$
-$d\geq 2$, $q$ odd $q^{2d^{2}+d}$ $q^{d}$ $\dfrac{1}{2d+1}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $q^{2d^{2}+d}$ $q^{2d-2}$ $\dfrac{2(d-1)}{d(2d+1)}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $q^{2d^{2}-d}$ $q^{2d-3}$ $\dfrac{2d-3}{d(2d-1)}$
-Table 3. Asymptotic quasirandom degree of simple classical groups of Lie type.
-$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments Asymptotic for $|L|$ Asymptotic for $\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{6}(q)$ $q^{78}$ $q^{11}$ $\dfrac{11}{78}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{7}(q)$ $q^{133}$ $q^{17}$ $\dfrac{17}{133}$
-$\dfrac{c1^{1}}{1_{1_{1}}}$ $E_{8}(q)$ $q^{248}$ $q^{29}$ $\dfrac{29}{248}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ even $q^{52}$ $q^{11}$ $\dfrac{11}{52}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $q^{52}$ $q^{8}$ $\dfrac{2}{13}$
+Table 2. $\mathbf{D}_{0}(L)$ for simple classical groups.
+In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups [7].
+We remark that, after $(\mathrm{PSL}_{2}(q))_{q\geq 4}$, which is asymptotically $\frac{1}{3}$-quasirandom, the most quasirandom families of finite simple groups are $(\mathrm{Sp}_{4}(q))_{q\,even}$ and the Suzuki groups $(^{2}B_{2}(q^{2}))_{q\geq 8}$, whose quasirandom degrees are asymptotically $\frac{3}{10}$.
+$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\sim\log_{q}|L|$ $\sim\log_{q}\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $2d^{2}+d$ $12d-$ 1 $\dfrac{2d-1}{2d^{2}+d}$
+$d\geq 2$, $q$ odd $2d^{2}+d$ $d$ $\dfrac{1}{2d+1}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $2d^{2}+d$ $22d-$ 2 $\dfrac{2d-2}{2d^{2}+d}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $2d^{2}-d$ $32d-$ 3 $\dfrac{2d-3}{2d^{2}-d}$
+Table 3. Asymptotic quasirandom degree of simple classical groups.
+$\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\sim\log_{q}|L|$ $\sim\log_{q}\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{6}(q)$ $78$ $11$ $\dfrac{11}{78}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $E_{7}(q)$ $133$ $17$ $\dfrac{17}{133}$
+$\dfrac{c1^{1}}{1_{1_{1}}}$ $E_{8}(q)$ $248$ $29$ $\dfrac{29}{248}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ even $52$ $11$ $\dfrac{11}{52}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $52$ $8$ $\dfrac{2}{13}$
 […]
 […]
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}E_{6}(q^{2})$ $q^{78}$ $q^{11}$ $\dfrac{11}{78}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{3}D_{4}(q^{3})$ $q^{28}$ $q^{5}$ $\dfrac{5}{28}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}F_{4}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $q^{52}$ $q^{11}$ $\dfrac{11}{52}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $q^{10}$ $q^{3}$ $\dfrac{3}{10}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $q^{14}$ $q^{4}$ $\dfrac{2}{7}$
-Table 4. Asymptotic quasirandom degree of simple exceptional groups of Lie type.
+$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}E_{6}(q^{2})$ $78$ $11$ $\dfrac{11}{78}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{3}D_{4}(q^{3})$ $28$ $5$ $\dfrac{5}{28}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}F_{4}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $52$ $11$ $\dfrac{11}{52}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $10$ $3$ $\dfrac{3}{10}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $14$ $4$ $\dfrac{2}{7}$
+Table 4. Asymptotic quasirandom degree of simple exceptional groups.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ $|L|$ $\mathbf{D}_{0}(L)$ $\bm{\varepsilon}(L)$ $L$ $|L|$ $\mathbf{D}_{0}(L)$ $\bm{\varepsilon}(L)$
@@
 Table 5. Quasirandom degree of the sporadic simple groups.
+Acknowledgments: We thank S. Eberhard for enlightening conversations on the paper, and P. Tiep for useful informations about his articles.

### v3

@@
 Quasirandom and Quasisimple Groups
-Abstract.
+Abstract
 Fix $\varepsilon>0$.
@@
 In this paper, we give a structure theorem for large $\varepsilon$-quasirandom groups, and we completely classify the $\frac{1}{5}$-quasirandom groups.
-Key words and phrases:
-Quasirandom groups, quasisimple groups.
-2020 Mathematics Subject Classification:
-Primary: 20D99, 20C99.
-1. Introduction
+[…]
+[…]
+[classification=text]
+1 Introduction
 Informally, a finite group is called quasirandom if its nontrivial complex irreducible representations have “large” degree, where the precise meaning of “large” depends on context.
@@
 Fix $\varepsilon>0$.
-If $G$ is an $\varepsilon$-quasirandom group, then $|G/Sol(G)|\geq|G|^{\varepsilon}$.
-If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/Sol(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
-Conversely, suppose that $G$ is perfect, $|G/Sol(G)|\geq|G|^{\varepsilon}$, and that $G/Sol(G)$ is $\varepsilon$-quasirandom.
+If $G$ is an $\varepsilon$-quasirandom group, then $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$.
+If $|G|$ is sufficiently large depending on $\varepsilon$, then $G/\mathrm{Sol}(G)$ is a direct product of boundedly many finite simple groups of Lie type of bounded rank.
+Conversely, suppose that $G$ is perfect, $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$, and that $G/\mathrm{Sol}(G)$ is $\varepsilon$-quasirandom.
 If $|G|$ is sufficiently large depending on $\varepsilon$, then $G$ is $\frac{\varepsilon^{2}}{3}$-quasirandom.
@@
 It is a widespread belief that $(\mathrm{PSL}_{2}(q))_{q\geq 4}$ is the “most quasirandom” family of groups [23, 28], but not much is known.
-Not only this is actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
+Not only is this actually the case, but we prove that the vast majority of highly quasirandom groups are quasisimple, shedding a good light on the issue.
 Theorem B.
@@
 quantitative) analysis of quasirandom degrees, modular and projective representations.
-The paper is structured as follows.
+The paper is organized as follows.
 In Section 2 we introduce the (exponential) quasirandom degree $\bm{\varepsilon}(G)$ of a finite group, and give basic results about quasisimple groups, modular and projective representations.
-In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem Bb.
+In Section 3, we provide some highly quasirandom nonquasisimple groups: they are of affine type as in Theorem B a-b.
 We prove the main theorems in Section 4.
-Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups, so we collect these data in the Appendix.
-2. Preliminaries
-2.1. Minimal degree of a nontrivial representation
+Fundamental to our discussion is the knowledge of the quasirandom degrees of the finite simple groups, so we collect these data in the Appendix A.
+2 Preliminaries
+2.1 Minimal degree of a nontrivial representation
 In this paper, every group is finite. For a nontrivial finite group $G$, we denote by $\mathbf{D}_{0}(G)$ the minimal degree of a nontrivial irreducible complex representation of $G$, that is,
 $\mathbf{D}_{0}(G)\>:=\>\min\left\{\dim(\pi)\mid\pi\in\mathrm{Irr}(G)\setminus\{1\}\right\}.$
+Observe that the veracity of Corollaries C and D implies that $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group $G$, and $\mathbf{D}_{0}(G)<|G|^{0.2143}$ if $G$ is not quasisimple.
 We start with the following, which appears as [25, Ex. 1.3.2].
@@
 Lev [16] proved that $\bm{\iota}(G)\leq|G|^{0.5}$ for every finite group $G$ whose order is not a prime (equality is achieved by groups of order $p^{2}$).
-From Corollaries C and D, we obtain $\mathbf{D}_{0}(G)<|G|^{0.2143}$ whenever $G$ is not quasisimple, and $\mathbf{D}_{0}(G)<|G|^{0.3464}$ for every finite group.
-2.2. Quasirandom degree
+2.2 Quasirandom degree
 Let $G$ be a nontrivial finite group. We introduce the (exponential) quasirandom degree of $G$ by
@@
 Let $G$ be a finite perfect group.
-The following statements are roughly equivalent, in the sense that if one statement holds for a small constant $\varepsilon>0$, then all others hold with constants that are a positive fraction of $\varepsilon$:
-  • $G$ is $\varepsilon$-quasirandom;
-  • every proper subgroup of $G$ has index at least $|G|^{\varepsilon}$;
-  • for every subset $S$ of size at least $|G|^{1-\varepsilon}$ we have $S^{3}=G$;
-  • every product-free subset of $G$ has size at most $|G|^{1-\varepsilon}$.
-Let $N\lhd G$ be a proper normal subgroup. From an easy computation, we write
+The following statements are roughly equivalent, in the sense that if one statement holds for a constant $\varepsilon_{i}>0$, then all others hold with constant $\varepsilon_{j}=c_{i,j}\varepsilon_{i}$, for some $c_{i,j}>0$.
+  • $G$ is $\varepsilon_{1}$-quasirandom;
+  • every proper subgroup of $G$ has index at least $|G|^{\varepsilon_{2}}$;
+  • for every subset $S$ of size at least $|G|^{1-\varepsilon_{3}}$ we have $S^{3}=G$;
+  • every product-free subset of $G$ has size at most $|G|^{1-\varepsilon_{4}}$.
+Let $G$ be a finite group and $N\lhd G$ a proper normal subgroup. We can write
 $\bm{\varepsilon}(G)\>=\>\bm{\varepsilon}(G/N)\cdot\frac{\log\mathbf{D}_{0}(G)}{\log\mathbf{D}_{0}(G/N)}\cdot\frac{\log|G/N|}{\log|G|}.$
@@
 Hence, substituting $|G|=|G/N||N|$, we obtain
-(2.1) $|N|\>\leq\>\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.$
-What makes (2.1) so useful is that in many occasions we barely know that $\bm{\varepsilon}(G)$ is large, but we have a good knowledge of $G/N$.
+$|N|\>\leq\>\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.$ (1)
+What makes (1) so useful is that in many occasions we barely know that $\bm{\varepsilon}(G)$ is large, but we have a good knowledge of $G/N$.
 The next lemma provides a gap between $\bm{\varepsilon}(G)$ and $\bm{\varepsilon}(G/N)$ when $N$ is abelian but noncentral, and it greatly simplifies the proof of Theorem B.
@@
 Proof.
-Since $N$ is abelian, we have $N\subseteq\mathbf{C}_{G}(N)\neq G$. Combining Lemma 2.1, Corollary 2.3(ii), and (2.1), we write
+Since $N$ is abelian, we have $N\subseteq\mathbf{C}_{G}(N)\neq G$. Combining Lemma 2.1, Corollary 2.3(ii), and (1), we write
 $\begin{split}\mathbf{D}_{0}(G/N)&\leq\mathbf{D}_{0}(G/\mathbf{C}_{G}(N))\\ &<|N|\\ &\leq\frac{\mathbf{D}_{0}(G/N)^{1/\bm{\varepsilon}(G)}}{|G/N|}.\end{split}\,$
@@
 Lemma 2.8.
-Let $A$ and $B$ be two finite groups, and suppose that $\bm{\varepsilon}(A)\leq\bm{\varepsilon}(B)$. Then
-$\frac{\bm{\varepsilon}(A)}{2}\>\leq\>\bm{\varepsilon}(A\times B)\>\leq\>\frac{\bm{\varepsilon}(B)}{2}.$
+The following holds:
+  • (i)
+    If $\bm{\varepsilon}(A)\leq\bm{\varepsilon}(B)$, then $\bm{\varepsilon}(A\times B)\leq\frac{\bm{\varepsilon}(B)}{2}$;
+  • (ii)
+    if $(G_{i})_{i=1}^{n}$ are finite groups, then $\bm{\varepsilon}\left(\prod_{i=1}^{n}G_{i}\right)<\frac{1}{n}$.
 Proof.
 We have $\mathbf{D}_{0}(A\times B)=\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}$, so that we can write
-$\begin{split}\bm{\varepsilon}(A\times B)&=\frac{\log\left(\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}\right)}{\log(|A||B|)}\\ &=\frac{\log\left(\min\{|A|^{\bm{\varepsilon}(A)},|B|^{\bm{\varepsilon}(B)}\}\right)}{\log|A|+\log|B|}\\ &\leq\bm{\varepsilon}(B)\cdot\frac{\min\{\log|A|,\log|B|\}}{\log|A|+\log|B|}\\ &\leq\frac{\bm{\varepsilon}(B)}{2}.\end{split}\,$
-The proof of the other inequality is similar. ∎
-Corollary 2.9.
-Let $(G_{i})_{i=1}^{n}$ be finite groups and suppose $\bm{\varepsilon}(G_{1})\leq...\leq\bm{\varepsilon}(G_{n})$.
-Then $\bm{\varepsilon}(A_{1})/n\leq\bm{\varepsilon}(\prod_{i=1}^{n}A_{i})\leq\bm{\varepsilon}(A_{n})/n$.
-2.3. Quasisimple groups
+$\begin{split}\bm{\varepsilon}(A\times B)&=\frac{\log\left(\min\{\mathbf{D}_{0}(A),\mathbf{D}_{0}(B)\}\right)}{\log(|A||B|)}\\ &=\frac{\log\left(\min\{|A|^{\bm{\varepsilon}(A)},|B|^{\bm{\varepsilon}(B)}\}\right)}{\log|A|+\log|B|}\\ &\leq\bm{\varepsilon}(B)\cdot\frac{\min\{\log|A|,\log|B|\}}{\log|A|+\log|B|}.\end{split}\,$
+Point (i) follows. For (ii), observe that there is a group, say $G_{n}$, satisfying $\log|G_{n}|\leq\frac{1}{n}\sum_{i=1}^{n}\log|G_{i}|$. It follows from the previous computation that
+$\bm{\varepsilon}\left(\prod_{i=1}^{n}G_{i}\right)<1\cdot\frac{\min\{\sum_{i=1}^{n-1}\log|G_{i}|,\log|G_{n}|\}}{\sum_{i=1}^{n}\log|G_{n}|}\leq\frac{1}{n}.$
+The proof is complete. ∎
+2.3 Quasisimple groups
 We recall that a central extension of a group $Q$ is a short exact sequence
@@
 33].
-The next two lemmas are essentially well known. We write down Lemma 2.11 below, which is a little more general, and will be useful later.
+The next two lemmas are essentially well known. We write down the proof of Lemma 2.10 below, which is a little more general, and will be useful later.
 Lemma 2.9.
@@
 By Grün’s Lemma [12] we have $\mathbf{Z}(G/Z)=1$, so $\mathbf{Z}(G/N)\subseteq Z/N$.
-Now from the hypotheses we have $N_{0}/N\subseteq Z/N$, i.e. $N_{0}\subseteq Z$. ∎
-We conclude this subsection with the following observation, which is useful for an inductive argument.
+Now from the hypotheses we have $N_{0}/N\subseteq Z/N$, i.e.
+$N_{0}\subseteq Z$.
+∎
+We conclude this subsection with the following observation, which is useful in an inductive argument.
 Lemma 2.11.
@@
 For each $i\in\{1,2\}$, we have
-$[h,g]N_{i}=[hN_{i},gN_{i}]\subseteq N_{i}.$
-It follows that $[h,g]\in N_{1}\cap N_{2}=1$, and since $h\in H$ is arbitrary, we conclude $H\subseteq\mathbf{Z}(G)$.
-Finally, by Lemma 2.10, $G$ itself is quasisimple, against our initial assumption.
-∎
-2.4. Linear and projective representations
+$h,g]N_{i}=[hN_{i},gN_{i}]\subseteq N_{i}.$
+It follows that $h,g]\in N_{1}\cap N_{2}=1$, and since $h\in H$ is arbitrary, we conclude $H\subseteq\mathbf{Z}(G)$. Finally, by Lemma 2.9, $G$ itself is quasisimple, against our initial assumption. ∎
+2.4 Linear and projective representations
 So far, we have only dealt with complex linear representations, but modular and projective representations will be a key tool in the proof of Theorem B.
@@
 Equivalently, it is a group homomorphism
-$\rho:\>G\to\mathrm{PGL}(V).$
+$\rho\colon\>G\to\mathrm{PGL}(V).$
 The degree of the representation is the dimension of $V$.
@@
 It follows immediately from the definition. ∎
-For the proof of Theorem B, we need detailed informations about the $\frac{1}{4}$-quasirandom finite simple groups. The infinite families of $\frac{1}{4}$-quasirandom simple groups are
+For the proof of Theorem B, we need detailed information about the $\frac{1}{4}$-quasirandom finite simple groups. The infinite families of $\frac{1}{4}$-quasirandom simple groups are
 $\mathrm{PSL}_{d}(q)\hbox{ with }d\in\{2,3\},\quad\mathrm{Sp}_{4}(q)\hbox{ with }q\hbox{ even},$
-$G_{2}(q)\hbox{ with }q=3(\mathrm{mod}\,6),\quad^{2}B_{2}(q^{2}),\quad^{2}G_{2}(q^{2})\,.$
-The relevant informations can be recovered from [15, 24] for classical groups, and from [17, Sec. 4] for exceptional groups (see also the Appendix). We collect these data in Table 1.
+$G_{2}(q)\hbox{ with }q=3\>(\mathrm{mod}\,6),\quad^{2}B_{2}(q^{2}),\quad^{2}G_{2}(q^{2})\,.$
+The relevant information can be recovered from [15, 24] for classical groups, and from [17, Sec. 4] for exceptional groups (see also the Appendix). We collect these data in Table 1.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\mathbf{D}_{0}(L)$ $\mathbf{R}_{0}(L)$ $\mathbf{R}_{r}(L)$, $r\neq p$ $\mathbf{R}_{p}(L)$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{2}(q)$ $q$ even $q-1$ $q-1$ $q-1$ $2$
-[…]
-[…]
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{3}(q)$ $q\neq 2,4$ $q(q+1)$ $q(q+1)$ $q^{2}+q-1$ $3$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{4}(q)$ $q$ even $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $4$
-[…]
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $4$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $\phi_{12}(q)$ $\phi_{12}(q)$ $\phi_{12}(q)-\delta_{2}$ $7$
-Table 1. The $\frac{1}{4}$-quasirandom simple groups of Lie type.
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{2}(q)$ $q$ even q1q-1 q1q-1 q1q-1 22
+q=1(mod 4)q=1\,($\mathrm${mod}\,4) $\dfrac{q+1}{2}$ $\dfrac{q+1}{2}$ $\dfrac{q-1}{2}$ 22
+q=3(mod 4)q=3\,($\mathrm${mod}\,4) $\dfrac{q-1}{2}$ $\dfrac{q-1}{2}$ $\dfrac{q-1}{2}$ 22
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{3}(q)$ $q\neq 2,4$ $q(q+1)$ $q(q+1)$ $q^{2}+q-1$ 33
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{Sp}_{4}(q)$ $q$ even $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ $\dfrac{1}{2}q(q-1)^{2}$ 44
+$\dfrac{1^{1}}{1_{1_{1}}}$ $G_{2}(q)$ q=3(mod 6)q=3($\mathrm${mod}\,6) $\phi_{3}(q)\phi_{6}(q)$ $\phi_{3}(q)\phi_{6}(q)$ $\phi_{3}(q)\phi_{6}(q)-\delta_{1}$ 77
+$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ $\dfrac{\sqrt{2}}{2}q\phi_{1}(q)\phi_{2}(q)$ 44
+$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $\phi_{12}(q)$ $\phi_{12}(q)$ $\phi_{12}(q)-\delta_{2}$ 77
+Table 1: The $\frac{1}{4}$-quasirandom simple groups of Lie type.
 We denote by $\phi_{\ell}(q)$ the $\ell$-th cyclotomic polynomial in the variable $q$.
-The value of $\delta_{1}$ is $1$ if $r=2$, $q=3^{f}$, $f\geq 2$, or $r=3$, $3$ divides $q-1$, and it is 00 otherwise.
-The value of $\delta_{2}$ is $1$ if $r=2$, and 00 otherwise.
+The value of $\delta_{1}$ is $1$ if $r=2$, $q=3^{f}$, $f\geq 2$, or $r=3$, 33 divides q1q-1, and it is 0 otherwise.
+The value of $\delta_{2}$ is $1$ if $r=2$, and 0 otherwise.
 Remark 2.13.
@@
 Moreover, we point out that the comparison between $\mathbf{R}_{0}(L)$ and $\mathbf{R}_{r}(L)$ ($r\neq p$) can be explicitly found, for all simple groups of Lie type, in [26, Sec. 3].
-3. Quasirandom groups of affine type
+3 Quasirandom groups of affine type
 In this short section, we provide some highly quasirandom nonquasisimple groups.
@@
 Before giving the results, we revise some notions from the representation theory of finite groups.
-Let $V$ be a finite vector space, and let $V^{*}$ be its dual space.
-We remark that $V^{*}\cong Hom(V,\mathbb{C}^{*})$.
-If $\phi:V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}:V^{*}\rightarrow V^{*}$ is defined by
-$\phi^{*}(f)\>:=\>\phi\circ f.$
-Let $K$ be a finite group. If $\rho:K\rightarrow\mathrm{GL}(V)$ is a representation of $K$, then the dual representation $\rho^{*}:K\rightarrow\mathrm{GL}(V^{*})$ is defined by
+Let $V$ be a vector space over $\mathbb{C}$, and let $V^{*}$ be its dual space.
+We remark that $V^{*}\cong\mathrm{Hom}(V,\mathbb{C}^{*})$.
+If $\phi\colon V\rightarrow V$ is a homomorphism of vector spaces, then $\phi^{*}\colon V^{*}\rightarrow V^{*}$ is defined by
+$\phi^{*}(f)\>:=\>f\circ\phi.$
+Let $K$ be a finite group. If $\rho\colon K\rightarrow\mathrm{GL}(V)$ is a representation of $K$, then the dual representation $\rho^{*}\colon K\rightarrow\mathrm{GL}(V^{*})$ is defined by
 $\rho^{*}(g)\>:=\>\rho(g^{-1})^{*}$
@@
 A direct computation shows that $\bm{\varepsilon}(H_{q})$ lies between $5^{-1}-q^{-1}$ and $5^{-1}$ for all $q$, so in particular $\lim_{q\rightarrow+\infty}\bm{\varepsilon}(H_{q})=\left(\frac{1}{5}\right)^{-}$.
-4. Proof of the main theorems
-4.1. Proof of Theorem A
+4 Proof of the main theorems
+4.1 Proof of Theorem A
 We immediately prove the second part of the theorem. We restate it as follows.
@@
 Fix $\varepsilon,\delta>0$, and let $G$ be a perfect group.
-Suppose that $G/Sol(G)$ is $\varepsilon$-quasirandom, and that $|G/Sol(G)|\geq|G|^{\delta}$.
+Suppose that $G/\mathrm{Sol}(G)$ is $\varepsilon$-quasirandom, and that $|G/\mathrm{Sol}(G)|\geq|G|^{\delta}$.
 If $|G|$ is sufficiently large with respect to $\varepsilon$ and $\delta$, then $G$ is $\frac{\varepsilon\delta}{3}$-quasirandom.
@@
 By [8, Lemma 4.10], we have
-$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/Sol(G))^{1/2},$
-where $c$ is an absolute constant. Looking back at [19], we check that $c=10^{-10}$ works. It follows that
-[…]
+$\mathbf{D}_{0}(G)\>\geq\>c\,\mathbf{D}_{0}(G/\mathrm{Sol}(G))^{1/2},$
+where $c$ is an absolute constant. It follows that
+$|G|^{\bm{\varepsilon}(G)}\>\geq\>c|G/\mathrm{Sol}(G)|^{\bm{\varepsilon}(G/\mathrm{Sol}(G))/2},$
 and passing to the logarithms,
-$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/Sol(G))}{2}\cdot\frac{\log|G/Sol(G)|}{\log|G|}-\frac{10\log 10}{\log|G|}.$
+$\bm{\varepsilon}(G)\>\geq\>\frac{\bm{\varepsilon}(G/\mathrm{Sol}(G))}{2}\cdot\frac{\log|G/\mathrm{Sol}(G)|}{\log|G|}-\frac{\log(c^{-1})}{\log|G|}.$
 The proof follows. ∎
 The main idea behind the other direction of Theorem A is the following.
-Under the hypothesis that $Sol(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$.
+Under the hypothesis that $\mathrm{Sol}(G)=1$, we prove that $G=F^{*}(G)$, the generalized Fitting subgroup of $G$.
 If this is not the case, then $G/F^{*}(G)$ has a nontrivial simple quotient $L$.
@@
 A similar argument will be used in a portion of the proof of Theorem B.
-Going into details, we start reporting the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
+Going into details, we start recording the Schreier conjecture, which is now a theorem as a consequence of the Classification of the Finite Simple Groups.
 Theorem 4.2 (Schreier conjecture).
@@
 Let $N\lhd G$, and suppose that a section $H_{0}/N_{0}\cong L$ of $G$ is simple.
-Then $N\subseteq N_{0}$, or $NN_{0}=H_{0}$.
+Then $N\subseteq N_{0}$, or $NN_{0}\geqslant H_{0}$.
 In the first case $L$ appears as a section of $G/N$, while in the second case $L$ appears as a section of $N$.
@@
 Fix $\varepsilon>0$.
-From Corollary 2.3(i), it is clear that $|G/Sol(G)|\geq|G|^{\varepsilon}$.
-Moreover, $G/Sol(G)$ is again $\varepsilon$-quasirandom.
-Up to replacing $G$ with $G/Sol(G)$, we can assume that $G$ has a trivial solvable radical.
+From Corollary 2.3(i), it is clear that $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$.
+Moreover, $G/\mathrm{Sol}(G)$ is again $\varepsilon$-quasirandom.
+Up to replacing $G$ with $G/\mathrm{Sol}(G)$, we can assume that $G$ has a trivial solvable radical.
 Now it is well known that the generalized Fitting subgroup $F=F^{*}(G)$ coincides with the socle of $G$, and it is the direct product of nonabelian simple groups $T_{i}$’s.
@@
 By the definition of quasirandom degree, we have
-(4.1) $|G|\>\leq\>\mathbf{D}_{0}(L)^{1/\varepsilon}.$
+$|G|\>\leq\>\mathbf{D}_{0}(L)^{1/\varepsilon}.$ (2)
 So, if $\mathbf{D}_{0}(L)$ is bounded, then $|G|$ is bounded.
@@
 From Lemma 2.2 we obtain $\mathbf{D}_{0}(L)\leq n$, and so
-[…]
-When $\mathbf{D}_{0}(L)$ is sufficiently large, this gives a contradiction to (4.1).
-We have proven that $G/Sol(G)=F^{*}(G/Sol(G))$, so $G/Sol(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/Sol(G)=\prod_{i}(T_{i})^{n_{i}}$.
-From Remark 4.5, the $T_{i}$’s are all simple groups of Lie type.
+$|F|\>\geq\>60^{\mathbf{D}_{0}(L)}.$
+When $\mathbf{D}_{0}(L)$ is sufficiently large, this gives a contradiction to (2).
+We have proven that $G/\mathrm{Sol}(G)=F^{*}(G/\mathrm{Sol}(G))$, so $G/\mathrm{Sol}(G)$ is an $\varepsilon$-quasirandom direct product of finite simple groups, say $G/\mathrm{Sol}(G)=\prod_{i}(T_{i})^{n_{i}}$.
+From Remark 4.5, each $T_{i}$ is a simple group of Lie type of $\varepsilon$-bounded rank.
 Let $N=\sum_{i}n_{i}$.
-From Corollary 2.9, we have $\varepsilon\leq\mathbf{D}_{0}(G/Sol(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$.
-Moreover, each block $(T_{i})^{n_{i}}$ has to be $\varepsilon$-quasirandom as a quotient of $G$.
-From Corollary 2.9 and Remark 4.5, we have
-$\varepsilon\>\leq\>\bm{\varepsilon}((T_{i})^{n_{i}})\>=\>\frac{\bm{\varepsilon}(T_{i})}{n_{i}}\>\leq\>\frac{10}{n_{i}\,rank(T_{i})},$
-so that $n_{i}\,rank(T_{i})\leq 10\varepsilon^{-1}$ for each $i$. ∎
+From Lemma 2.8(ii), we have $\varepsilon\leq\bm{\varepsilon}(G/\mathrm{Sol}(G))\leq N^{-1}$, so that $N\leq\varepsilon^{-1}$.
+∎
 Remark 4.6.
-An attentive reader will have noticed that a bound of the type $\mathbf{D}_{0}(G)\geq c\log|G|$ is strong enough to give that $G/Sol(G)$ is the direct product of nonabelian finite simple groups.
-It is natural to ask whether the condition $|G/Sol(G)|\geq|G|^{\varepsilon}$ is actually superflous in the second part of Theorem A, as it may be implied by the quasirandomness of the sole $G/Sol(G)$.
-The answer is no, as it can even happen that $|G/Sol(G)|$ is bounded: it is enough to consider the perfect deleted permutation module $(C_{n})^{4}\rtimes\mathrm{Alt}(5)$ when $n$ is large.
-The following example, where $|G/Sol(G)|$ is unbounded but too small, is due to S.
+An attentive reader will have noticed that a bound of the type $\mathbf{D}_{0}(G)\geq c\log|G|$ is strong enough to give that $G/\mathrm{Sol}(G)$ is the direct product of nonabelian finite simple groups.
+It is natural to ask whether the condition $|G/\mathrm{Sol}(G)|\geq|G|^{\varepsilon}$ is actually superflous in the second part of Theorem A, as it may be implied by the quasirandomness of the sole $G/\mathrm{Sol}(G)$.
+The answer is no, as it can even happen that $|G/\mathrm{Sol}(G)|$ is bounded: it is enough to consider the perfect deleted permutation module $(C_{n})^{4}\rtimes\mathrm{Alt}(5)$ when $n$ is large.
+The following example, where $|G/\mathrm{Sol}(G)|$ is unbounded but too small, is due to S.
 Eberhard.
@@
 It can be proven that $G_{d}$ is perfect, and its solvable radical is the subgroup where $g\in\mathbf{Z}(\mathrm{SL}_{2}(p))$ is a scalar matrix.
-In particular, $G_{d}/Sol(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $0.34$-quasirandom for large $p$, but its cardinality is less than $p^{3}$.
-4.2. Proof of Theorem B
-We first prove the shallow version of Theorem B. This is a nice application of (2.1).
+In particular, $G_{d}/\mathrm{Sol}(G_{d})$ is isomorphic to $\mathrm{PSL}_{2}(p)$, which is $0.34$-quasirandom for large $p$, but its cardinality is less than $p^{3}$.
+4.2 Proof of Theorem B
+We first prove the shallow version of Theorem B. This is a nice application of (1).
 Proof of Corollary D.
@@
 We have $\bm{\varepsilon}(G/N)\geq\frac{1}{3}$, and $G/N$ is isomorphic to either $J_{1}$ or $O^{\prime}N$.
-In the former case, from (2.1), we have
-[…]
-So $N=1$, a contradiction. Now suppose $G/N\cong O^{\prime}N$. From (2.1),
-[…]
+In the former case, from (1), we have
+$|N|\leq\frac{\mathbf{D}_{0}(J_{1})^{1/\bm{\varepsilon}(G)}}{|J_{1}|}\leq\frac{56^{3}}{175\,560}\approx 1.00031.$
+So $N=1$, a contradiction. Now suppose $G/N\cong O^{\prime}N$. From (1),
+$|N|\leq\frac{\mathbf{D}_{0}(O^{\prime}N)^{1/\bm{\varepsilon}(G)}}{|O^{\prime}N^{\prime}|}\leq\frac{10\,944^{3}}{460\,815\,505\,920}\approx 2.84446,$
 so that $N$ is the cyclic group of order $2$.
@@
 $G/(M_{1}\cap M_{2})=M_{1}/(M_{1}\cap M_{2})\times M_{2}/(M_{1}\cap M_{2})\cong G/M_{2}\times G/M_{1}.$
-Lemma 2.8 yields
+Lemma 2.8(i) yields
 $\bm{\varepsilon}(G)\leq\bm{\varepsilon}((G/M_{2})\times(G/M_{1}))<\frac{1}{5},$
 and $\bm{\varepsilon}(G)<\frac{1}{6}$ if neither $J_{1}$ or $O^{\prime}N$ appears as a quotient of $G$. If one among $J_{1}$ or $O^{\prime}N$ appears as a quotient, then
-[…]
+$|G|\leq\mathbf{D}_{0}(G)^{6}<\mathbf{D}_{0}(O^{\prime}N)^{6}=10\,944^{6}<2\cdot 10^{24},$
 because $\mathbf{D}_{0}(J_{1})<\mathbf{D}_{0}(O^{\prime}N)$. ∎
@@
 The key point is that the (complex, linear) quasirandomness of $L$ controls also the degrees of the modular projective representations.
-This gives a lower bound on $|N|$, which we put against the upper bound given by (2.1).
+This gives a lower bound on $|N|$, which we put against the upper bound given by (1).
 We start with the details.
@@
     $G/N$ is a perfect central extension of $(\mathbb{F}_{q})^{4}\rtimes\mathrm{Sp}_{4}(q)$, $q=2^{f}$, $f\geq 6$.
-If we are in case $(a)$, using (2.1) we obtain
-[…]
-Hence $N$ is trivial, a contradiction.
+If we are in case $(a)$, using (1) we obtain
+$|N|\leq\frac{\mathbf{D}_{0}((\mathbb{F}_{3})^{3}\rtimes\mathrm{SL}_{3}(3))^{5}}{|(\mathbb{F}_{3})^{3}\rtimes\mathrm{SL}_{3}(3)|}=\frac{12^{5}}{151\,632}\approx 1,64.$
+Hence NN is trivial, a contradiction.
 If we are in case $(b)$, we remark that $G/M\cong\mathrm{Sp}_{4}(q)$ and $\bm{\varepsilon}(G/N)<\frac{3}{14}$.
@@
 We have
-(4.2) $|G|\>\leq\>\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$
+$|G|\>\leq\>\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}.$ (3)
 Considering the embedding of $G$ into $\mathrm{Aut}(T^{n})$, and arguing as before, we obtain $\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))\leq n$.
-Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (4.2).
+Since $|N|\geq 60^{n}$, and $60^{x}>x^{5}$ for every $x\geq 1$, this contradicts (3).
 $G/N$ is quasisimple. Let $C=\mathbf{C}_{G}(N)$. We distinguish three cases.
@@
 Moreover, the action of $G/N$ on $N$ by conjugation provides a projective modular representation of $L$ (recall that $G/N$ is quasisimple).
-As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (2.1).
+As already explained, the key idea is that the quasirandomness of $G/N$ gives a lower bound on $|N|$, which we put against the upper bound given by (1).
 We observe that $\mathrm{Alt}(5)$ and $\mathrm{Alt}(6)$ are $\frac{1}{4}$-quasirandom, but if one among them appears as $L$, then
-[…]
+$|G|\leq\mathbf{D}_{0}(L)^{5}\leq 5^{5}=3\,125,$
 and we can easily work wit a computer (see Remark 4.9). One last time, we divide the discussion into three cases:
@@
     $L$ is a sporadic simple group.
-Let $L$ be as described in $(i)$. We remark that $|L|>\mathbf{D}_{0}(L)^{3}$ by Corollary D. Hence, by (2.1),
+Let $L$ be as described in $(i)$. We remark that $|L|>\mathbf{D}_{0}(L)^{3}$ by Corollary D. Hence, by (1),
 $|N|\leq|M|\leq\frac{\mathbf{D}_{0}(L)^{5}}{|L|}<\mathbf{D}_{0}(L)^{2}.$
@@
 The above inequality is possible only if $\mathbf{D}_{0}(L)\leq 6$, which gives
-[…]
+$|G|\leq\mathbf{D}_{0}(L)^{5}\leq 6^{5}=7\,776.$
 This case can be handled computationally.
-Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (2.1), we find
+Let $L$ be as described in $(ii)$. We first deal with $L\cong\mathrm{Sp}_{4}(q)$ with $q$ even, $q\geq 4$. Since the Schur multiplier is trivial, we have $N=M$. By (1), we find
 $\begin{split}|N|&\leq\frac{\mathbf{D}_{0}(\mathrm{Sp}_{4}(q))^{5}}{|\mathrm{Sp}_{4}(q)|}\\ &\leq\frac{(q^{2}-1)^{5}(q^{2}-q)^{5}}{2^{5}(q+1)^{5}}\cdot\frac{1}{q^{4}(q^{4}-1)(q^{2}-1)}\\ &<q^{5}.\end{split}$
@@
 So $|N|=q^{4}$, and we are in the situation of Theorem Bb.
-The argument to deal with the other cases is the same, using (2.1), Lemma 2.13 and Table 1. When $L\cong\mathrm{PSL}_{2}(q)$,
+The argument to deal with the other cases is the same, using (1), Lemma 2.12 and Table 1. When $L\cong\mathrm{PSL}_{2}(q)$,
 $|N|\leq\frac{\mathbf{D}_{0}(\mathrm{PSL}_{2}(q))^{5}}{|\mathrm{PSL}_{2}(q)|}\leq\frac{(q-1)^{5}}{q(q^{2}-1)}<q^{2}.$
@@
 $q^{3}\leq|N|\leq\frac{q^{5}(q+1)^{5}}{q^{3}(q^{3}-1)(q^{2}-1)(q-1)}<q^{3}.$
-[…]
+When $L\cong G_{2}(q)$, q=3(mod 6)q=3($\mathrm${mod}\,6), we have
 $q^{7}\leq|N|\leq\frac{(q+1)^{5}(q^{2}-q+1)^{5}}{q^{6}(q^{6}-1)(q^{2}-1)}<q^{7}.$
@@
 Therefore $N=M$ is the unique normal subgroup of $G$, and $N$ can be thought as the vector space of a nontrivial modular representation of $L$.
-For each sporadic group, we use [14] to obtain a lower bound on $|N|$, and we compare the result with (2.1).
+For each sporadic group, we use [14] to obtain a lower bound on $|N|$, and we compare the result with (1).
 For $L\cong M_{11}$, we have
@@
 For $L\cong J_{1}$,
-[…]
+$2^{20}\leq|N|\leq\frac{56^{5}}{|J_{1}|}\leq 3\,138;$
 For $L\cong J_{3}$,
@@
 Finally, for $L\cong O^{\prime}N$,
-[…]
+$3^{154}\leq|N|\leq\frac{10\,944^{5}}{|O^{\prime}N|}\leq 10^{10}.$
 These contradictions conclude the proof of Theorem B.
-Appendix: Quasirandom degree of finite simple groups
+Appendix
+Appendix A Appendix: Quasirandom degree of finite simple groups
 We already discussed the alternating groups in Remark 4.5. In particular, we observed that they are the less quasirandom among the nonabelian finite simple groups.
@@
 See also [18].
-In Table 2, we use these informations to record $\mathbf{D}_{0}(L)$ for the simple classical groups.
+In Table 2, we use this information to record $\mathbf{D}_{0}(L)$ for the simple classical groups.
 We remark that $\mathbf{R}_{0}(L)<\mathbf{D}_{0}(L)$ holds in some cases, for example for some symplectic groups in odd characteristic.
+In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups [7].
+We remark that, after $(\mathrm{PSL}_{2}(q))_{q\geq 4}$, which is asymptotically $\frac{1}{3}$-quasirandom, the most quasirandom families of finite simple groups are $(\mathrm{Sp}_{4}(q))_{q=2^{f}}$ and the Suzuki groups $(^{2}B_{2}(q^{2}))_{q\geq 8}$, whose quasirandom degrees are asymptotically $\frac{3}{10}$.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\mathbf{D}_{0}(L)$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d=2$, $q$ even $q-1$
-[…]
-[…]
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d=2$, $q$ even q1q-1
+$d=2$, q=1(mod 4)q=1\,($\mathrm${mod}\,4) $\dfrac{q+1}{2}$
+$d=2$, q=3(mod 4)q=3\,($\mathrm${mod}\,4) $\dfrac{q-1}{2}$
 $d\geq 3$, $(d,q)\neq(3,2),(4,2),(4,3)$ $\dfrac{q^{d}-q}{q-1}$
-$d=3$, $q=2$ $3$
-$d=4$, $q=2$ $7$
+$d=3$, $q=2$ 33
+$d=4$, $q=2$ 77
 $d=4$, $q=3$ $26$
@@
 $d\geq 4$ even, $(d,q)\neq(4,2)$ $\dfrac{q^{d}-1}{q+1}\leq\mathbf{D}_{0}(L)\leq\dfrac{q^{d}+q}{q+1}$
-$d=4$, $q=2$ $5$
+$d=4$, $q=2$ 55
 $\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
-[…]
-[…]
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q\neq 3$ and $12(q^{2}-1)\nmid q^{2d}-$ 1 $\dfrac{q^{2d}-1}{q^{2}-1}$
-$d\geq 3$, $q=3$ or $12(q^{2}-1)\mid q^{2d}-$ 1 $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
+$d\geq 2$, $q^{d}=1\,(\mathrm{mod}\,4)$ $\dfrac{q^{d}+1}{2}$
+$d\geq 2$, $q^{d}=3\,(\mathrm{mod}\,4)$ $\dfrac{q^{d}-1}{2}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q\neq 3$ and $2(q^{2}-1)\nmid q^{2d}-1$ $\dfrac{q^{2d}-1}{q^{2}-1}$
+$d\geq 3$, $q=3$ or $2(q^{2}-1)\mid q^{2d}-1$ $\dfrac{(q^{d}-1)(q^{d}-q)}{2(q+1)}$
 $\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{-}_{2d}(q)$ $d\geq 4$ $\dfrac{(q^{d}+1)(q^{d-1}-q)}{q^{2}-1}$
@@
 $d\geq 5$, $q=2$ or $d\geq 4$, $q=3$ $\dfrac{(q^{d}-1)(q^{d-1}-1)}{q^{2}-1}$
-Table 2. $\mathbf{D}_{0}(L)$ for simple classical groups.
-In Tables 3, 4, and 5, we collect the asymptotic quasirandom degrees of the finite simple groups of Lie type, and the quasirandom degrees of the sporadic simple groups [7].
-We remark that, after $(\mathrm{PSL}_{2}(q))_{q\geq 4}$, which is asymptotically $\frac{1}{3}$-quasirandom, the most quasirandom families of finite simple groups are $(\mathrm{Sp}_{4}(q))_{q\,even}$ and the Suzuki groups $(^{2}B_{2}(q^{2}))_{q\geq 8}$, whose quasirandom degrees are asymptotically $\frac{3}{10}$.
+Table 2: $\mathbf{D}_{0}(L)$ for simple classical groups.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\sim\log_{q}|L|$ $\sim\log_{q}\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $d^{2}-1$ $d-1$ $\dfrac{1}{d+1}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $2d^{2}+d$ $12d-$ 1 $\dfrac{2d-1}{2d^{2}+d}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSL}_{d}(q)$ $d\geq 2$ $d^{2}-1$ d1d-1 $\dfrac{1}{d+1}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSU}_{d}(q)$ $d\geq 3$ $d^{2}-1$ d1d-1 $\dfrac{1}{d+1}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{PSp}_{2d}(q)$ $d\geq 2$, $q$ even $2d^{2}+d$ 2d12d-1 $\dfrac{2d-1}{2d^{2}+d}$
 $d\geq 2$, $q$ odd $2d^{2}+d$ $d$ $\dfrac{1}{2d+1}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $2d^{2}+d$ $22d-$ 2 $\dfrac{2d-2}{2d^{2}+d}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $2d^{2}-d$ $32d-$ 3 $\dfrac{2d-3}{2d^{2}-d}$
-Table 3. Asymptotic quasirandom degree of simple classical groups.
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\Omega_{2d+1}(q)$ $d\geq 3$, $q$ odd $2d^{2}+d$ 2d22d-2 $\dfrac{2d-2}{2d^{2}+d}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $\mathrm{P}\Omega^{\pm}_{2d}(q)$ $d\geq 4$ $2d^{2}-d$ 2d32d-3 $\dfrac{2d-3}{2d^{2}-d}$
+Table 3: Asymptotic quasirandom degree of simple classical groups.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ Comments $\sim\log_{q}|L|$ $\sim\log_{q}\mathbf{D}_{0}(L)$ $\lim\limits_{q\rightarrow+\infty}\bm{\varepsilon}(L)$
@@
 $\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ even $52$ $11$ $\dfrac{11}{52}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $52$ $8$ $\dfrac{2}{13}$
-[…]
-[…]
+$\dfrac{1^{1}}{1_{1_{1}}}$ $F_{4}(q)$ $q$ odd $52$ 88 $\dfrac{2}{13}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $G_{2}(q)$ $q\neq 3\,(\mathrm{mod}\,6)$ $14$ 33 $\dfrac{3}{14}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $G_{2}(q)$ q=3(mod 6)q=3\,($\mathrm${mod}\,6) $14$ 44 $\dfrac{2}{7}$
 $\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}E_{6}(q^{2})$ $78$ $11$ $\dfrac{11}{78}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{3}D_{4}(q^{3})$ $28$ $5$ $\dfrac{5}{28}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $^{3}D_{4}(q^{3})$ $28$ 55 $\dfrac{5}{28}$
 $\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}F_{4}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $52$ $11$ $\dfrac{11}{52}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $10$ $3$ $\dfrac{3}{10}$
-$\dfrac{1^{1}}{1_{1_{1}}}$ ${}^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $14$ $4$ $\dfrac{2}{7}$
-Table 4. Asymptotic quasirandom degree of simple exceptional groups.
+$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}B_{2}(q^{2})$ $q=\sqrt{2^{2m+1}}$ $10$ 33 $\dfrac{3}{10}$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $^{2}G_{2}(q^{2})$ $q=\sqrt{3^{2m+1}}$ $14$ 44 $\dfrac{2}{7}$
+Table 4: Asymptotic quasirandom degree of simple exceptional groups.
 $\dfrac{1^{1}}{1_{1_{1}}}$ $L$ $|L|$ $\mathbf{D}_{0}(L)$ $\bm{\varepsilon}(L)$ $L$ $|L|$ $\mathbf{D}_{0}(L)$ $\bm{\varepsilon}(L)$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{11}$ 7 92079207\,9207 920 $10$ $\approx 0.2565$ $Fi_{24}$ 125 520 $125\,520\,570\cdot 10^{16}$ 8 67186718\,6718 671 $\approx 0.1635$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{12}$ 95 0409504095\,04095 040 $11$ $\approx 0.2093$ $HS$ 44 352 0004435200044\,352\,00044 352 000 $22$ $\approx 0.1756$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{22}$ 443 520443520443\,520443 520 $21$ $\approx 0.2342$ $McL$ 898 128 000898128000898\,128\,000898 128 000 $22$ $\approx 0.1500$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{23}$ 10 200 9601020096010\,200\,96010 200 960 $22$ $\approx 0.1916$ $He$ 4 030 387 20040303872004\,030\,387\,2004 030 387 200 $51$ $\approx 0.1778$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{24}$ 244 823 040244823040244\,823\,040244 823 040 $23$ $\approx 0.1624$ $Ru$ 145 926 $145\,926\,144\cdot 10^{3}$ $378$ $\approx 0.2309$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{1}$ 175 560175560175\,560175 560 $56$ $\approx 0.3334$ $Suz$ 448 345 $448\,345\,497\cdot 10^{3}$ $143$ $\approx 0.1850$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{2}$ 604 800604800604\,800604 800 $14$ $\approx 0.1924$ $O^{\prime}N$ 460 815 $460\,815\,505\cdot 10^{3}$ 10 9441094410\,94410 944 $\approx 0.3464$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{3}$ 50 232 9605023296050\,232\,96050 232 960 $85$ $\approx 0.2506$ $HN$ 273 030 $273\,030\,912\cdot 10^{6}$ $133$ $\approx 0.1472$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{4}$ 867 755 $867\,755\,710\cdot 10^{11}$ 1 33313331\,3331 333 $\approx 0.1568$ $Ly$ 517 651 $517\,651\,790\cdot 10^{8}$ 2 48024802\,4802 480 $\approx 0.2031$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{3}$ 495 766 $495\,766\,656\cdot 10^{3}$ $23$ $\approx 0.1165$ $Th$ 907 459 $907\,459\,438\cdot 10^{8}$ $248$ $\approx 0.1413$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{2}$ 423 054 $423\,054\,213\cdot 10^{5}$ $23$ $\approx 0.1000$ $B$ 415 478 $415\,478\,148\cdot 10^{24}$ 4 37143714\,3714 371 $\approx 0.1083$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{1}$ 415 777 $415\,777\,680\cdot 10^{10}$ $276$ $\approx 0.1311$ $M$ 808 017 $808\,017\,424\cdot 10^{44}$ 196 883196883196\,883196 883 $\approx 0.0983$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{22}$ 645 617 $645\,617\,516\cdot 10^{5}$ $78$ $\approx 0.1371$ ${}^{2}F_{4}(2)^{\prime}$ 17 971 2001797120017\,971\,20017 971 200 $104$ $\approx 0.2781$
-$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{23}$ 408 947 $408\,947\,047\cdot 10^{10}$ $782$ $\approx 0.1555$
-Table 5. Quasirandom degree of the sporadic simple groups.
-Acknowledgments: We thank S. Eberhard for enlightening conversations on the paper, and P. Tiep for useful informations about his articles.
+$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{11}$ 7 9207\,920 $10$ $\approx 0.2565$ $Fi_{24}$ $125\,520\,570\cdot 10^{16}$ 8 6718\,671 $\approx 0.1635$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{12}$ 95 04095\,040 $11$ $\approx 0.2093$ HSHS 44 352 00044\,352\,000 $22$ $\approx 0.1756$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{22}$ 443 520443\,520 2121 $\approx 0.2342$ McLMcL 898 128 000898\,128\,000 $22$ $\approx 0.1500$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{23}$ 10 200 96010\,200\,960 $22$ $\approx 0.1916$ HeHe 4 030 387 2004\,030\,387\,200 $51$ $\approx 0.1778$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $M_{24}$ 244 823 040244\,823\,040 $23$ $\approx 0.1624$ RuRu $145\,926\,144\cdot 10^{3}$ $378$ $\approx 0.2309$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{1}$ 175 560175\,560 $56$ $\approx 0.3334$ SuzSuz $448\,345\,497\cdot 10^{3}$ $143$ $\approx 0.1850$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{2}$ 604 800604\,800 $14$ $\approx 0.1924$ $O^{\prime}N$ $460\,815\,505\cdot 10^{3}$ 10 94410\,944 $\approx 0.3464$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{3}$ 50 232 96050\,232\,960 $85$ $\approx 0.2506$ HNHN $273\,030\,912\cdot 10^{6}$ $133$ $\approx 0.1472$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $J_{4}$ $867\,755\,710\cdot 10^{11}$ 1 3331\,333 $\approx 0.1568$ LyLy $517\,651\,790\cdot 10^{8}$ 2 4802\,480 $\approx 0.2031$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{3}$ $495\,766\,656\cdot 10^{3}$ $23$ $\approx 0.1165$ ThTh $907\,459\,438\cdot 10^{8}$ $248$ $\approx 0.1413$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{2}$ $423\,054\,213\cdot 10^{5}$ $23$ $\approx 0.1000$ $B$ $415\,478\,148\cdot 10^{24}$ 4 3714\,371 $\approx 0.1083$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $Co_{1}$ $415\,777\,680\cdot 10^{10}$ $276$ $\approx 0.1311$ $M$ $808\,017\,424\cdot 10^{44}$ 196 883196\,883 $\approx 0.0983$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{22}$ $645\,617\,516\cdot 10^{5}$ $78$ $\approx 0.1371$ $^{2}F_{4}(2)^{\prime}$ 17 971 20017\,971\,200 $104$ $\approx 0.2781$
+$\dfrac{1^{1}}{1_{1_{1}}}$ $Fi_{23}$ $408\,947\,047\cdot 10^{10}$ $782$ $\approx 0.1555$
+Table 5: Quasirandom degree of the sporadic simple groups.
+Acknowledgments
+We thank an anonymous referee who found a mistake in Lemma 2.8. Moreover, we thank S. Eberhard and P.H. Tiep for useful conversations.